### v1

# Eugene Lerman

Abstract.
We introduce Poisson $C^{\infty}$-rings and Poisson local $C^{\infty}$-ringed spaces. We show that the spectrum of a Poisson $C^{\infty}$-ring is an affine Poisson $C^{\infty}$-scheme. We then discuss applications that include singular symplectic and Poisson reductions, singular quasi-Poisson reduction, coisotropic reduction and Poisson-Dirac subschemes.
1. Introduction
In early 1970s Meyer [M] and, independently, Marsden and Weinstein [MW] developed a mathematical framework that connected symmetries of Hamiltonian systems with reduction of the number of degrees of freedom. Their work lead to a flurry of activity that persisted for several decades. It was quickly observed that in many interesting cases the Meyer-Marsden-Weinstein reduced spaces (a.k.a. symplectic quotients) are singular. Unlike algebraic geometry where singular varieties and schemes make perfectly good sense, a singular manifold is an oxymoron. While singular spaces arose in differential geometry long before symplectic reduction and a wide variety of ways of dealing with singularities has been developed (which I will not attempt to survey), there is no consensus about the “best” approach.
In the case of singular symplectic quotients the approaches roughly fall into three categories: “geometric” “algebraic” (cf. [AGJ]) and “homological” [St]. In this paper we connect geometric and algebraic approaches using tools of algebraic geometry over $C^{\infty}$-rings [Du, J]. (The connections with homological approach should come out of derived differential geometry once it is well developed.) From the $C^{\infty}$-ring point of view the algebraic approach results in Poisson $C^{\infty}$-rings (a concept that we introduce in this paper) while geometric approach results in spaces with a structure sheaf of Poisson $C^{\infty}$-rings. We show that Dubuc’s spectrum functor from the category of $C^{\infty}$-rings to the category of local $C^{\infty}$-ringed spaces send Poisson $C^{\infty}$-rings to Poisson ringed spaces (see Theorem 3.18). This unifies algebraic and geometric views of reduction in symplectic and Poisson geometry. It also makes singular reduction a lot simpler since there is no need to worry about stratifications. With almost no extra work we also makes sense of singular quasi-Poisson reduction as Poisson $C^{\infty}$-schemes. Other applications include coisotropic reduction in Poisson geometry and Poisson-Dirac submanifolds and more generally subschemes.
Recall that a $C^{\infty}$-ring is a product preserving functor from the category of coordinate vector spaces $\mathbb{R}^{n}$, $n\in\mathbb{N}$, and smooth maps to the category of sets [MR, J]. Equivalently a $C^{\infty}$-ring is a set $\mathscr{A}$ together with a collection of $n$-ary operations, $n\geq 0$, one $n$-ary operation
$f_{\mathscr{A}}:\mathscr{A}^{n}\to\mathscr{A}$
for each smooth function $f\in C^{\infty}(\mathbb{R}^{n})$. These operations are associative. Maps between $C^{\infty}$-rings are natural transformation between functors. Equivalently they are maps of sets preserving the operations. It follows quickly from the definition that any $C^{\infty}$-ring has an underlying $\mathbb{R}$-algebra. A $C^{\infty}$-ring derivation $v:\mathscr{A}\to\mathscr{A}$ of a $C^{\infty}$-ring $\mathscr{A}$ is an $\mathbb{R}$-linear map that satisfies a generalization of the Leibniz rule. Namely, for any $n$, for any $f\in C^{\infty}(\mathbb{R}^{n})$ and for any $a_{1},\ldots,a_{n}\in\mathscr{A}$ one requires that
(1.1) $v(f_{\mathscr{A}}(a_{1},\ldots,a_{n}))=\sum_{i=1}^{n}(\partial_{i}f)_{\mathscr{A}}(a_{1},\ldots,a_{n})\cdot v(a_{i}).$
For example, if $M$ is a manifold then its algebra $C^{\infty}(M)$ of smooth functions is a $C^{\infty}$-ring with the operations are given by compositions. Namely, for all $n$, all $f\in C^{\infty}(\mathbb{R}^{n})$ and all $a_{1},\ldots,a_{n}\in C^{\infty}(M)=C^{\infty}(M,\mathbb{R})$
$f_{C^{\infty}(M)}(a_{1},\ldots,a_{n}):=f\circ(a_{1},\ldots,a_{n}).$
And then a vector field $v$ on a manifold $M$ is a $C^{\infty}$-ring derivation since
$v(f\circ(a_{1},\ldots,a_{n}))=\sum_{i=1}^{n}\big{(}\partial_{i}f\circ(a_{1},\ldots,a_{n})\big{)}\cdot v(a_{i}).$
Recall that a Poisson algebra $A$ over a commutative ring $k$ is an associative (and often commutative) $k$-algebra with a compatible Lie algebra bracket {,}$\{\cdot,\cdot\}:A\times A\to A$ such that for each $a\in A$ the map $\{a,\cdot\}:A\to A$ is a $k$-algebra derivation. Its is well-known [Po, K1] that the algebra-geometric spectrum of a Poisson algebra is a Poisson scheme, that is, a local ringed space whose structure sheaf is a sheaf of Poisson algebras. Poisson schemes have proven to be useful in algebraic geometry and representation theory. However, if one starts with a Poisson manifold $M$ and regards its algebra of functions $C^{\infty}(M)$ as Poisson algebra over $\mathbb{R}$ then the associated Poisson scheme is not the original manifold with its sheaf of smooth functions. Of course one may feel that there is no need to view Poisson manifolds as affine Poisson schemes, and it’s hard to argue with this point of view. On the other hand the Sniatycki-Weinstein algebraic reduction [SW] produces a Poisson algebra over $\mathbb{R}$, which is a Poisson algebra of a manifold under a regularity assumption. However, in general this Poisson algebra need not be an algebra of functions on any space [AGJ]. It is natural to wonder if this algebra is the algebra of global sections of a sheaf on some space. This question was the original motivation for undertaking the work in this paper.
Additionally, as was mentioned above, in symplectic and Poisson geometry there often arise geometric objects that are singular but are not algebraic varieties or schemes in any natural way. Examples include singular symplectic quotients [ACG, SL], singular Poisson quotients, closures of coadjoint orbits of real Lie groups and, more generally, closures of symplectic leaves in Poisson manifolds, and so on. The techniques developed in this paper give a novel point of view of such objects — see Section 4.
Organization of the paper
In Section 2 we define Poisson $C^{\infty}$-rings. We show that if a $C^{\infty}$-ring $\mathscr{A}$ is Poisson and $I\subset\mathscr{A}$ is a Poisson ideal then the quotient $\mathscr{A}/I$ is again a Poisson $C^{\infty}$-ring. The point here that the induced bracket on $\mathscr{A}/I$ is a $C^{\infty}$-ring biderivation, and not just an $\mathbb{R}$-algebra biderivation. We then give examples of such quotients and relate one of them to the “universal” reduction of Arms, Cushman and Gotay [ACG]. We also show that if $\mathscr{A}$ is a Poisson $C^{\infty}$-ring and $I\subset\mathscr{A}$ is a multiplicative ideal then its Poisson normalizer $N(I)$ is a $C^{\infty}$-subring of $\mathscr{A}$.
In Section 3 we review Joyce’s construction of Dubuc’s spectrum functor
$\operatorname{Spec}:{C^{\infty}\mathsf{Ring}}^{\mbox{\sf{\tiny{op}}}}\to\mathsf{LC^{\infty}RS}$
($C^{\infty}\mathsf{Ring}$ is the category of $C^{\infty}$-rings, $\mathsf{LC^{\infty}RS}$ is the category of local $C^{\infty}$-ringed spaces) and prove that if $\mathscr{A}$ is a Poisson $C^{\infty}$-ring then the affine $C^{\infty}$-scheme $\operatorname{Spec}(\mathscr{A})$ is Poisson (i.e., its structure sheaf is a sheaf of Poisson $C^{\infty}$-rings). We also prove that if $\varphi:\mathscr{A}\to\mathscr{B}$ is a map of Poisson $C^{\infty}$-rings then $\operatorname{Spec}(\varphi):\operatorname{Spec}(\mathscr{B})\to\operatorname{Spec}(\mathscr{A})$ is Poisson.
In Section 4 we discuss various applications of Theorem 3.18: symplectic reduction, Poisson reduction, quasi-Poisson reduction, coisotropic reduction in the setting of Poisson manifolds, Snyaticki-Weinstein reduction as a Poisson $C^{\infty}$-scheme and Poisson-Dirac submanifolds and subschemes.
There are two appendices. In Appendix A we address the question of getting one’s hands on $\operatorname{Spec}(\mathscr{A})$. We show that if a $C^{\infty}$-ring $\mathscr{A}$ is finitely generated (as a $C^{\infty}$-ring!) then the space $X_{\mathscr{A}}$ of its points embeds as a closed subspace in some $\mathbb{R}^{N}$ (the embedding depends on the choice of generators). If $G\times M\to M$ is an action of a Lie group on a manifold $M$ then the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ of invariant functions may or may not be finitely generated (if $G$ is compact and the number of orbit types is finite then $C^{\infty}(M)^{G}$ is finitely generated by a nontrivial theorem of Schwarz [Sch]). Regardless of $C^{\infty}(M)^{G}$ being finitely generated we show that for proper actions (of not necessarily compact) Lie groups the space $X_{C^{\infty}(M)^{G}}$ of points the ring $C^{\infty}(M)^{G}$ of invariants is homeomorphic to the orbit space $M/G$ and consequently $\operatorname{Spec}(C^{\infty}(M)^{G})$ is isomorphic to $(M/G,C^{\infty}_{M/G})$ where $C^{\infty}_{M/G}\subset C^{0}_{M/G}$ is a sheaf of “smooth” functions on the orbit space. This implies, in particular, if $M$ is Poisson, the action of $G$ on $M$ is Hamiltonian, $\mu:M\to\mathfrak{g}^{\vee}$ is an equivariant moment map, $\alpha\in\mathfrak{g}^{\vee}$ and $I_{\alpha}\subset C^{\infty}(M)^{G}$ is the vanishing ideal of the level set $\mu^{-1}(\alpha)$ then the set of points $X_{C^{\infty}(M)^{G}/I_{\alpha}}$ embeds into the orbit space $M/G$ as a closed subspace and the $C^{\infty}$-Poisson scheme $\operatorname{Spec}(C^{\infty}(M)^{G}/I_{\alpha})$ “is” a closed subscheme of $(M/G,C^{\infty}_{M/G})$. The same result, mutatis mutandis holds if $M$ is a quasi-Poisson manifold.
In Appendix B we discuss the Cushman-Weinstein conjecture [Eg1, Eg2, SL] on embedding of symplectic quotients into Euclidean spaces.
2. Poisson $C^{\infty}$-rings
We start by introducing notation, which is not entirely standard. It is, however, convenient.
Notation 2.1.
$C^{\infty}$-rings form a category $C^{\infty}\mathsf{Ring}$. The objects of this category are $C^{\infty}$-rings and morphisms are functions between $C^{\infty}$-rings that preserve all the operations. That is, $\varphi:\mathscr{A}\to\mathscr{B}$ is a map of $C^{\infty}$-rings if for every $n\geq 0$, all $f\in C^{\infty}(\mathbb{R}^{n})$ and all $a_{1},\ldots,a_{n}\in\mathscr{A}$
$\varphi(f_{\mathscr{A}}(a_{1},\ldots,a_{n})=f_{\mathscr{B}}(\varphi(a_{1}),\ldots,\varphi(a_{n})).$
Here, as in the introduction, $f_{\mathscr{A}}:\mathscr{A}^{n}\to\mathscr{A}$ is the operation defined by the function $f$.
Similarly a $C^{\infty}$-ringed space is a pair $(X,\mathscr{O}_{X})$ where $X$ is a topological space and $\mathscr{O}_{X}$ is a sheaf of $C^{\infty}$-rings. It is a local $C^{\infty}$-ringed space if all the stalks of the sheaf $\mathscr{O}_{X}$ are local $C^{\infty}$-rings (see Definition 3.1 below). A map/morphism $\underline{f}:(X,\mathscr{O}_{X})\to a(Y,\mathscr{O}_{Y})$ from a local $C^{\infty}$-ringed space $(X,\mathscr{O}_{X})$ to a local $C^{\infty}$-ringed space $(Y,\mathscr{O}_{Y})$ is a pair of maps $(f,f_{\#})$ where $f:X\to Y$ is a continuous map and $f_{\#}:\mathscr{O}_{Y}\to f_{*}\mathscr{O}_{\mathsf{X}}$ is a map of sheaves of $C^{\infty}$-rings. Note that unlike ordinary algebraic geometry the map on stalks induced by $f_{\#}$ automatically preserves the unique maximal ideals.
Notation 2.2.
We denote the category of local $C^{\infty}$-ringed spaces and their maps by $\mathsf{LC^{\infty}RS}$.
Definition 2.3.
A module over a $C^{\infty}$-ring $\mathscr{A}$ is a module over the underlying $\mathbb{R}$-algebra (cf. [J, L1]).
Definition 2.4.
Let $\mathscr{M}$ be a module over a $C^{\infty}$-ring $\mathscr{A}$. A $C^{\infty}$-ring derivation (or a $C^{\infty}$-derivation for short) with values in the module $\mathscr{M}$ is an $\mathbb{R}$-linear map $v:\mathscr{A}\to\mathscr{M}$ so that for any $n$, any $f\in C^{\infty}(\mathbb{R}^{n})$ and any $a_{1},\ldots,a_{n}\in\mathscr{A}$
$v\left(f_{\mathscr{A}}(a_{1},\ldots,a_{n})\right)=\sum_{i=1}^{n}(\partial_{i}f)_{\mathscr{A}}(a_{1},\ldots,a_{n})\cdot v(a_{i}).$
Definition 2.5.
[…]
$ad(a)\equiv\{a,\cdot\}:\mathscr{A}\to\mathscr{A},\quad b\mapsto\{a,b\}$
is a $C^{\infty}$-ring derivation: for any $n>0$, any $f\in C^{\infty}(\mathbb{R}^{n})$ and any $a_{1},\ldots,a_{n}\in\mathscr{A}$
(2.1) $\left\{a,f_{\mathscr{A}}(a_{1},\ldots,a_{n})\right\}=\sum_{i=1}^{n}(\partial_{i}f)_{\mathscr{A}}(a_{1},\ldots,a_{n})\cdot\{a,a_{i}\}.$
Example 2.6.
Let $(P,\pi\in\Gamma\Lambda^{2}TM)$ be a Poisson manifold and $\{f,g\}=\langle df\wedge dg,\pi\rangle$ the corresponding bracket. Then $(C^{\infty}(P),\{\cdot,\cdot\})$ is a $C^{\infty}$-Poisson ring. This is because $ad(f)$ is a vector field for every $f\in C^{\infty}(P)$.
Definition 2.7.
A map of Poisson $C^{\infty}$-rings is a map of $C^{\infty}$-rings that preserve the Poisson brackets.
We define subalgebras of Poisson $C^{\infty}$-rings as in the case of ordinary Poisson algebras mutatis mutandis:
Definition 2.8.
A Poisson $C^{\infty}$-subring (i.e., a Poisson subalgebra) of a Poisson $C^{\infty}$-ring $(\mathscr{A},\{\cdot,\cdot\})$ is a $C^{\infty}$-subring $\mathscr{B}$ of $\mathscr{A}$ which is closed under the Poisson brackets.
Example 2.9.
Let $(M,\pi)$ be a Poisson manifold with an action of a Lie group $G$ that preserves the Poisson tensor $\pi$. Then the algebra $C^{\infty}(M)^{G}$ of $G$-invariant functions is a $C^{\infty}$-subring of $C^{\infty}(M)$ and a Poisson $C^{\infty}$-subring of $(C^{\infty}(M),\{\cdot,\cdot\})$.
Recall that an ideal $I$ in a $C^{\infty}$-ring $\mathscr{A}$ is an ordinary ideal in the $\mathbb{R}$-algebra underlying $\mathscr{A}$. Recall also that if $I$ is an ideal in the $C^{\infty}$-ring $\mathscr{A}$ then the quotient $\mathscr{A}/I$ is naturally a $C^{\infty}$-ring with the operations
$f_{\mathscr{A}/I}:(\mathscr{A}/I)^{n}\to\mathscr{A}/I$
given by
$f_{\mathscr{A}/I}(a_{1}+I,\ldots,a_{n}+I)=f_{\mathscr{A}}(a_{1},\ldots,a_{n})+I$
(for all $n\geq 0$, all $f\in C^{\infty}(\mathbb{R}^{n})$ and all $a_{1}+I,\ldots,a_{n}+I\in\mathscr{A}/I$). The fact that the operations are well-defined is a consequence of Hadamard’s lemma; see [MR].
Definition 2.10.
A Poisson ideal of a Poisson $C^{\infty}$-ring $\mathscr{A}$ is an ideal in (the $\mathbb{R}$-algebra underlying) $\mathscr{A}$ so that
$\{f,a\}\in I\qquad\textrm{ for all }f\in I,a\in\mathscr{A}.$
Definition 2.11.
A sheaf $\mathscr{O}$ of $C^{\infty}$-rings on a space $X$ is a sheaf of Poisson $C^{\infty}$-rings if for every open set $U\subset X$ the $C^{\infty}$-ring $\mathscr{O}(U)$ is Poisson and the restriction maps are maps of Poisson $C^{\infty}$-rings.
Lemma 2.12.
Let $I$ be a Poisson ideal in a Poisson $C^{\infty}$-ring $\mathscr{A}$. Then the quotient $C^{\infty}$-ring $\mathscr{A}/I$ is Poisson and the projection map $p:\mathscr{A}\to\mathscr{A}/I$ is map of Poisson $C^{\infty}$-rings.
Proof.
This is an easy consequence of Lemma 2.13 below. ∎
Lemma 2.13.
Let $v:\mathscr{A}\to\mathscr{A}$ be a $C^{\infty}$-ring derivation of a $C^{\infty}$-ring $\mathscr{A}$ and $I\subseteq\mathscr{A}$ an ideal with $v(I)\subset I$. Then the induced map
$\bar{v}:\mathscr{A}/I\to\mathscr{A}/I,\qquad\bar{v}(a+I):=v(a)+I$
is a $C^{\infty}$-ring derivation.
Proof.
Given $n$, $f\in C^{\infty}(\mathbb{R}^{n})$ and $a_{1},\ldots,a_{n}\in\mathscr{A}$ we compute:
$\begin{split}\bar{v}(f_{\mathscr{A}/I}(a_{1}+I,\ldots,a_{n}+I)=&\bar{v}(f_{\mathscr{A}}(a_{1},\ldots,a_{n})+I)\\ =&v(f_{\mathscr{A}}(a_{1},\ldots,a_{n}))+I\\ =&\left(\sum(\partial_{i}f)_{\mathscr{A}}(a_{1},\ldots,a_{n})v(a_{i})\right)+I\\ =&\sum((\partial_{i}f)_{\mathscr{A}}(a_{1},\ldots,a_{n})+I)(v(a_{i})+I)\\ =&\sum(\partial_{i}f)_{\mathscr{A}/I}(a_{1}+I,\ldots,a_{n}+I)\bar{v}(a_{i}+I).\end{split}$
The next lemma is useful for producing natural examples of Poisson ideals. It is a generalization of [BL, Lemma 2], which, in turn, is a variation on the ideas in [ACG].
Lemma 2.14.
Let $M$ be a Poisson manifold, and ${{\mathcal{A}}}$ be a Poisson $C^{\infty}$-subring of $C^{\infty}(M)$. Suppose that the Hamiltonian flows of functions in ${{\mathcal{A}}}$ preserve a subset $X$ of the manifold $M$. Then the vanishing ideal
${{\mathcal{I}}}(X):=\{f\in{{\mathcal{A}}}:f|_{X}=0\}$
of $X$ is a Poisson ideal of ${{\mathcal{A}}}$.
Proof.
Let $f$ be in ${{\mathcal{A}}}$, $x$ be a point in $X$ and $h$ be in the ideal ${{\mathcal{I}}}(X)$. Let $\gamma(t)$ denote the integral curve of the Hamiltonian vector field of $f$ with $\gamma(0)=x$. Then $\gamma(t)$ is in $X$ and consequently $h(\gamma(t))=0$ for all $t$. Differentiation with respect to $t$ yields
$0=\left.\frac{d}{dt}\right|_{0}h(\gamma(t))=\{f,h\}(x)$
Thus $\{f,h\}|_{X}=0$, i.e., $\{f,h\}$ is in the ideal ${{\mathcal{I}}}(X)$. ∎
Example 2.15.
Let $\mathfrak{g}$ be a real Lie algebra. Then its dual $\mathfrak{g}^{\vee}$ is canonically a Poisson manifold. Let $\mathscr{O}\subset\mathfrak{g}^{\vee}$ be a coadjoint orbit. Since the flows of Hamiltonian vector fields on $\mathfrak{g}^{\vee}$ preserve the orbit $\mathscr{O}$, the ideal of functions $I=\{f\in C^{\infty}(\mathfrak{g}^{\vee})\mid f|_{\mathscr{O}}=0\}$ is a Poisson ideal by Lemma 2.14. By Lemma 2.12 the $C^{\infty}$-ring $C^{\infty}(\mathfrak{g}^{\vee})/I$ is a Poisson $C^{\infty}$-ring. It’s the ring of (Whitney) smooth functions on the closure $\overline{\mathscr{O}}$ of the orbit $\mathscr{O}$: $C^{\infty}(\mathfrak{g}^{\vee})/I=C^{\infty}(\overline{\mathscr{O}}$). It follows that $C^{\infty}(\overline{\mathscr{O}})$ is a Poisson $C^{\infty}$-ring.
We can easily generalize the example above.
Example 2.16.
Let $(P,\pi)$ be a Poisson manifold and $X\subset P$ a closed subset which is a union of symplectic leaves of the Poisson tensor $\pi$. Then the ideal $I_{X}\subset C^{\infty}(P)$ of functions that vanish on $X$ is a Poisson ideal in the $C^{\infty}$-Poisson algebra $C^{\infty}(P)$.
Another application of Lemma 2.14 gives us the so called universal reduction of Arms, Cushman and Gotay [ACG]:
Example 2.17.
Let $(M,\omega)$ be a symplectic manifold with a Hamiltonian action of a Lie group $G$ and let $\mu:M\to\mathfrak{g}^{\vee}$ denote the corresponding equivariant moment map. Let $\alpha\in\mathfrak{g}^{\vee}$ be a point. Consider the set
$I_{\alpha}:=\{f\in C^{\infty}(M)^{G}\mid f|_{\mu^{-1}(\alpha)}=0\},$
the set of all invariant functions that vanish on the level set $\mu^{-1}(\alpha)$. Since the moment map $\mu$ is constant along the integral curves of all invariant functions $f\in C^{\infty}(M)^{G}$, the flows of invariant functions preserve $\mu^{-1}(\alpha)$. By Lemma 2.14 the ideal $I_{\alpha}$ is a Poisson ideal in the Poisson $C^{\infty}$-ring $cin(M)^{G}$. Therefore the quotient $C^{\infty}$-ring $C^{\infty}(M)^{G}/I_{\alpha}$ is a Poisson $C^{\infty}$-ring.
Remark 2.18.
Any element of the quotient $C^{\infty}$-ring $C^{\infty}(M)^{G}/I_{\alpha}$ gives rise to a well-defined continuous function on the quotient space $\mu^{-1}(\alpha)/G_{\alpha}$ where $G_{\alpha}$ is the stabilizer of $\alpha\in\mathfrak{g}^{\vee}$ (under the coadjoint action). In the next section (Theorem 3.18) we will see that one can associate to any Poisson $C^{\infty}$-ring $\mathscr{A}$ a space $X_{\mathscr{A}}$ with a sheaf of Poisson $C^{\infty}$-rings $\mathscr{O}_{\mathscr{A}}$. If the action of $G$ on $M$ is nice enough, the two spaces turn out to be homeomorphic. We will prove it in Appendix A.
Definition 2.19.
Let $(\mathscr{A},\{\cdot,\cdot\})$ be a Poisson $C^{\infty}$-ring. The normalizer $N(S)$ of a subset $S\subset\mathscr{A}$ is the set
$N(S):=\{a\in\mathscr{A}\mid\{s,a\}\in S\textrm{ for all }s\in S\}.$
Lemma 2.20.
Let $(\mathscr{A},\{\cdot,\cdot\})$ be a Poisson $C^{\infty}$-ring, $\mathscr{I}\subset\mathscr{A}$ an ideal in the commutative $\mathbb{R}$-algebra underlying $\mathscr{A}$. The normalizer $N(\mathscr{I})$ of $\mathscr{I}$ is a $C^{\infty}$-subring of $\mathscr{A}$. Consequently
$\mathscr{B}:=N(\mathscr{I})/(\mathscr{I}\cap N(\mathscr{I}))$
is a Poisson $C^{\infty}$-ring.
Proof.
Given $n>0$, $h\in C^{\infty}(\mathbb{R}^{n})$, $f_{1},\ldots f_{n}\in N(\mathscr{I})$ and $j\in\mathscr{I}$ we want to show that $\{j,h_{\mathscr{A}}(f_{1},\ldots f_{n})\}\in N(\mathscr{I})$. Since $\{j,\cdot\}:\mathscr{A}\to\mathscr{A}$ is a $C^{\infty}$-ring derivation,
$\{j,h_{\mathscr{A}}(f_{1},\ldots f_{n})\}=\sum_{i=1}^{n}(\partial_{i}h)_{\mathscr{A}}(f_{1},\ldots f_{n})\{j,f_{i}\}.$
Since $\mathscr{I}$ is an ideal and since $\{j,f_{i}\}\in\mathscr{I}$ for all $i$, it follows that $\{j,h_{\mathscr{A}}(f_{1},\ldots f_{n})\}\in N(\mathscr{I})$ as desired. ∎
Example 2.21 (Sniatycki-Weinstein algebraic reduction [SW]).
[…]
$\mathscr{B}=(C^{\infty}(M)/\mathscr{I})^{G}$
of $G$-invariant elements of the quotient $C^{\infty}(M)/\mathscr{I}$. They then argue [SW, Lemma 2] that $\mathscr{B}\simeq N(\mathscr{I})/\mathscr{I}$, where as in Lemma 2.20, $N(\mathscr{I})$ is the normalizer of the ideal $\mathscr{I}$ in the Poisson algebra $C^{\infty}(M)$. By Lemma 2.20, the Poisson algebra $\mathscr{B}$ is a Poisson $C^{\infty}$-ring.
Sniatycki and Weinstein (op. cit.) give an example of their reduced Poisson algebra that is not the Poisson algebra of a manifold. Arms, Gotay and Jennings [AGJ] give further examples where the reduced Poisson algebra is not an algebra of functions on a space.
As was noted in Remark 2.18 one can attach to the Sniatycki-Weinstein reduced algebra $\mathscr{B}=N(\mathscr{I})/\mathscr{I}$ (which is a Poisson $C^{\infty}$-ring) a space with a sheaf of Poisson $C^{\infty}$-rings. See Example 4.2 for further discussion.
3. Poisson $C^{\infty}$-schemes
The main result of this section is Theorem 3.18 and its relative version, Lemma 3.23: the spectrum of a Poisson $C^{\infty}$-ring is a Poisson $C^{\infty}$-scheme and if $\varphi:\mathscr{A}\to\mathscr{B}$ is a map of Poisson $C^{\infty}$-rings then $\operatorname{Spec}(\varphi):\operatorname{Spec}(\mathscr{B})\to\operatorname{Spec}(\mathscr{A})$ is a map of Poisson $C^{\infty}$-schemes. The proofs use the details of Joyce’s construction [J] of Dubuc’s spectrum functor $\operatorname{Spec}$ from the category ${C^{\infty}\mathsf{Ring}}^{\mbox{\sf{\tiny{op}}}}$ (the category opposite to the category of $C^{\infty}$-rings) to local $C^{\infty}$-ringed spaces $\mathsf{LC^{\infty}RS}$ (which will be defined below). Note that unlike the corresponding functor in algebraic geometry, Dubuc’s
$\operatorname{Spec}:{C^{\infty}\mathsf{Ring}}^{\mbox{\sf{\tiny{op}}}}\to\mathsf{LC^{\infty}RS}$
is not fully faithful. We start by reviewing some definitions that are used in Joyce’s construction.
Definition 3.1.
An $\mathbb{R}$-point of a $C^{\infty}$-ring $\mathscr{A}$ is a map of $C^{\infty}$-rings $p:\mathscr{A}\to\mathbb{R}$.
A $C^{\infty}$-ring $\mathscr{A}$ is local if it has a unique $\mathbb{R}$-point. Equivalently there exists a unique maximal ideal $M\subseteq\mathscr{A}$ so that the quotient $\mathscr{A}/M$ is isomorphic to $\mathbb{R}$.
A $C^{\infty}$-ringed space $(X,\mathscr{O}_{X})$ is a local $C^{\infty}$-ringed space if all the stalks of the structure sheaf $\mathscr{O}_{X}$ are local $C^{\infty}$-rings.
By analogy with algebraic geometry we define a Poisson local $C^{\infty}$-ringed space as follows.
Definition 3.2.
A Poisson local $C^{\infty}$-ringed space is a local $C^{\infty}$-ringed space $(X,\mathscr{O}_{X})$ so that the structure sheaf $\mathscr{O}_{X}$ is a sheaf of Poisson $C^{\infty}$-rings (cf. Definition 2.11).
Notation 3.3.
We denote the set of all $\mathbb{R}$-points of a $C^{\infty}$-ring $\mathscr{A}$ by $X_{\mathscr{A}}$. Thus
$X_{\mathscr{A}}:=\{p:\mathscr{A}\to\mathbb{R}\mid p\textrm{ is a map of $C^{\infty}$-rings}\}\equiv\operatorname{Hom}(\mathscr{A},\mathbb{R}).$
Remark 3.4.
It is well-known that if $M$ is a (second-countable Hausdorff) manifold then the map
(3.1) $M\to\operatorname{Hom}(C^{\infty}(M),\mathbb{R}),\qquad p\mapsto\operatorname{\mathsf{ev}}_{p}$
is a bijection (here and below $\operatorname{\mathsf{ev}}_{p}(f)=f(p)$ for all $f\in C^{\infty}(M)$). This is a theorem of Pursell [Pu, Chapter 8]. It is often referred to as Milnor’s exercise.
The surjectivity of (3.1) follows from the fact that $C^{\infty}(M)$ contains a proper function combined with Lemma A.7. The injectivity of (3.1) is the fact that smooth functions on Hausdorff manifolds separate points.
Construction 3.5 (The Zariski topology $\mathscr{T}_{\mathscr{A}}$ on the set $X_{\mathscr{A}}$ of $\mathbb{R}$-points of a $C^{\infty}$-ring $\mathscr{A}$).
The set $X_{\mathscr{A}}$ of $\mathbb{R}$-points of a $C^{\infty}$-ring $\mathscr{A}$ comes equipped with a natural topology, the Zariski topology. It is defined as follows: for $a\in\mathscr{A}$ let
$U_{a}:=\{p\in X_{\mathscr{A}}\mid p(a)\not=0\}.$
Since for a point $p$ of $\mathscr{A}$ and $a,b\in\mathscr{A}$
$p(ab)\not=0\quad\Leftrightarrow\quad p(a)\not=0\textrm{ and }p(b)\not=0,$
we have $U_{ab}=U_{a}\cap U_{b}$. Hence the collection $\{U_{a}\}_{a\in\mathscr{A}}$ of sets is a basis for a topology on $X_{\mathscr{A}}$, which we denote by $\mathscr{T}_{\mathscr{A}}$.
Remark 3.6.
By a theorem of Whitney any closed subset $C$ of a smooth (Hausdorff and paracompact) manifold $M$ is the set of zeros of some smooth function $f\in C^{\infty}(M)$. It follows that the bijection (3.1) is a homeomorphism between the manifold $M$ with its given topology and the set $X_{C^{\infty}(M)}$ of $\mathbb{R}$-points of $C^{\infty}(M)$ with the Zariski topology $\mathscr{T}_{C^{\infty}(M)}$.
Remark 3.7.
Unlike the Zariski topology in algebraic geometry the topology $\mathscr{T}_{\mathscr{A}}$ of Construction 3.5 is much nicer. It is Hausdorff and, additionally, it is completely regular: for any closed set $C\subset X_{\mathscr{A}}$ and for any point $x\not\in C$ there is a continuous function $f:X_{\mathscr{A}}\to\mathbb{R}$ with $f(x)\not=0$ and $f|_{C}=0$, see [J].
To construct the functor $\operatorname{Spec}$ on objects, given a $C^{\infty}$-ring $\mathscr{A}$ we need to equip the topological space $X_{\mathscr{A}}$ of its points with a sheaf $\mathscr{O}_{\mathscr{A}}$ of $C^{\infty}$-rings. To construct this structure sheaf $\mathscr{O}_{\mathscr{A}}$ we need to recall a few things about localizations of $C^{\infty}$-rings.
Lemma 3.8.
Given a $C^{\infty}$-ring $\mathscr{A}$ and a set $\Sigma$ of nonzero elements of $\mathscr{A}$ there exists a $C^{\infty}$-ring $\mathscr{A}\{\Sigma^{-1}\}$ and a map $\gamma:\mathscr{A}\to\mathscr{A}\{\Sigma^{-1}\}$ of $C^{\infty}$-rings with the following universal property:
  1. (3.8.i)
    $\gamma(a)$ is invertible in $\mathscr{A}\{\Sigma^{-1}\}$ for all $a\in\Sigma$ (i.e., there is $b\in\mathscr{A}\{\Sigma^{-1}\}$ such that $\gamma(a)b=1$) and
  2. (3.8.ii)
    for any map $\varphi:\mathscr{A}\to\mathscr{B}$ of $C^{\infty}$-rings so that $\varphi(a)$ is invertible in $\mathscr{B}$ for all $a\in\Sigma$ there exists a unique map $\overline{\varphi}:\mathscr{A}\{\Sigma^{-1}\}\to\mathscr{B}$ making the diagram
    ${\mathscr{A}}$${\mathscr{B}}$${\mathscr{A}\{\Sigma^{-1}\}}$$\scriptstyle{\varphi}$$\scriptstyle{\gamma}$$\scriptstyle{\overline{\varphi}}$
    commute.
Proof.
See [MR, p. 23]. ∎
Definition 3.9.
We refer to the map $\gamma:\mathscr{A}\to\mathscr{A}\{\Sigma^{-1}\}$ of Lemma 3.8 as a localization of the $C^{\infty}$-ring $\mathscr{A}$ at the set $\Sigma$.
Remark 3.10.
A localization of a $C^{\infty}$-ring $\mathscr{A}$ at a set $\Sigma$ is unique up to a unique isomorphism, so we can speak about the localization of $\mathscr{A}$ at $\Sigma$.
Notation 3.11.
Let $\mathscr{A}$ be a $C^{\infty}$-ring. By Lemma 3.8, for an $\mathbb{R}$-point $x:\mathscr{A}\to\mathbb{R}$ of a $C^{\infty}$-ring $\mathscr{A}$ there exists a localization of $\mathscr{A}$ at the set
$\{x\not=0\}:=\{a\in\mathscr{A}\mid x(a)\not=0\}$
We set
$\mathscr{A}_{x}:=\mathscr{A}\{\{x\not=0\}^{-1}\}.$
and denote the corresponding localization map by
(3.2) $\pi_{x}:\mathscr{A}\to\mathscr{A}_{x}$
Joyce proves [J, Proposition 2.14] that $\pi_{x}:\mathscr{A}\to\mathscr{A}_{x}$ is surjective with $I_{x}:=\ker\pi_{x}$ given by
(3.3) $I_{x}:=\{a\in\mathscr{A}\mid\textrm{ there is }d\in\mathscr{A}\textrm{ so that }x(d)\not=0\textrm{ and }ad=0\}.$
We think of $\mathscr{A}_{x}$ as the ring of germs of elements of $\mathscr{A}$ at the point $x$.
Remark 3.12.
In case of $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})$ and $x:C^{\infty}(\mathbb{R}^{n})\to\mathbb{R}$ (which is the evaluation at some point $p\in\mathbb{R}^{n}$ by Remark 3.4) the localization $(C^{\infty}(\mathbb{R}^{n}))_{x}$ is isomorphic to the usual ring of germs of functions at $p$. This is because both rings are localizations of $C^{\infty}(\mathbb{R}^{n})$ at the same set; see [J, Example 2.15].
Remark 3.13.
The localizations $\mathscr{A}_{x}$ are local rings. This is easy to see. Note first that for any $a\in\{x\not=0\}$, $x(a)$ is invertible in $\mathbb{R}$, hence $x:\mathscr{A}\to\mathbb{R}$ gives rise to $\overline{x}:\mathscr{A}_{x}\to\mathbb{R}$ with $x=\overline{x}\circ\pi_{x}$. Moreover, for any $c\in\mathscr{A}$, $\pi_{x}(c)\not\in\ker{\overline{x}}$ if and only if $x(c)\not=0$ if and only if $\pi_{x}(c)$ is invertible in $\mathscr{A}_{x}$. Hence $\mathscr{A}_{x}\smallsetminus\ker{\overline{x}}$ consists of units of $\mathscr{A}_{x}$ and therefore $\ker{\overline{x}}$ is a unique maximal ideal in $\mathscr{A}_{x}$.
Construction 3.14.
Given a $C^{\infty}$-ring $\mathscr{A}$ we construct the corresponding affine $C^{\infty}$-scheme $\operatorname{Spec}(\mathscr{A})$. In Construction 3.5 we defined a topology on the set $X_{\mathscr{A}}$ of $\mathbb{R}$-points of $\mathscr{A}$. We now construct the structure sheaf $\mathscr{O}_{\mathscr{A}}$ on $X_{\mathscr{A}}$.
We start by defining a candidate etale space $S_{\mathscr{A}}$ of the sheaf $\mathscr{O}_{\mathscr{A}}$:
$S_{\mathscr{A}}:=\bigsqcup_{x\in X_{\mathscr{A}}}\mathscr{A}_{x}\equiv\bigsqcup_{x\in X_{\mathscr{A}}}\mathscr{A}/I_{x}.$
The set $S_{\mathscr{A}}$ comes with the evident surjective map $\varpi:S_{\mathscr{A}}\to X_{\mathscr{A}}$ defined by $\varpi(s)=x$ for all $s\in\mathscr{A}_{x}\hookrightarrow S_{\mathscr{A}}$. For any $a\in\mathscr{A}$ we get a section $\mathfrak{s}_{a}:X_{\mathscr{A}}\to S_{\mathscr{A}}$ of $\varpi$:
(3.4) $\mathfrak{s}_{a}(x)=\pi_{x}(a)\equiv a_{x},$
where, as before, $\pi_{x}:\mathscr{A}\to\mathscr{A}_{x}$ is the localization map (3.2). The collection of sets
$\{\mathfrak{s}_{a}(U)\mid a\in\mathscr{A},U\in\mathsf{Open}(X_{\mathscr{A}})\}$
forms a basis for a topology on $S_{\mathscr{A}}$. In this topology the projection $\varpi:S_{\mathscr{A}}\to X_{\mathscr{A}}$ is a local homeomorphism and all sections $\mathfrak{s}_{a}:X_{\mathscr{A}}\to S_{\mathscr{A}}$ are continuous.
We define the structure sheaf $\mathscr{O}_{\mathscr{A}}$ of $\operatorname{Spec}(\mathscr{A})$ to be the sheaf of continuous sections of $\varpi:S_{\mathscr{A}}\to X_{\mathscr{A}}$. Equivalently
(3.5) $\begin{split}\mathscr{O}_{\mathscr{A}}(U)=&\{s:U\to\bigsqcup_{x\in U}\mathscr{A}_{x}\mid\textrm{ there is an open cover }\{U_{\alpha}\}_{\alpha\in A}\textrm{ of }U\\ &\textrm{ and }\{a_{\alpha}\}_{\alpha\in A}\subset\mathscr{A}\textrm{ so that }s|_{U_{\alpha}}=\mathfrak{s}_{a_{\alpha}}|_{U_{\alpha}}\textrm{ for all }\alpha\in A\}.\end{split}$
for every open subset $U$ of $X_{\mathscr{A}}$. The $C^{\infty}$-ring structure on the sets $\mathscr{O}_{\mathscr{A}}(U)$ is defined pointwise.
It follows from (3.5) that the sheaf $\mathscr{O}_{\mathscr{A}}$ is a sheafification of a presheaf ${\mathcal{P}}_{\mathscr{A}}$ defined by
(3.6) ${\mathcal{P}}_{\mathscr{A}}(U):=\{\mathfrak{s}_{a}|_{U}\mid a\in\mathscr{A}\}.$
It turns out that the stalk of the sheaf $\mathscr{O}_{\mathscr{A}}$ at a point $x\in X_{\mathscr{A}}$ is (isomorphic to) $\mathscr{A}_{x}$ ([J, Lemma 4.18]). Consequently the pair $(X_{\mathscr{A}},\mathscr{O}_{\mathscr{A}})$ is a local $C^{\infty}$-ringed space.
Remark 3.15.
Note that the canonical map $\eta:{\mathcal{P}}_{\mathscr{A}}\to\mathscr{O}_{\mathscr{A}}$ from a presheaf to its sheafification is simply the inclusion: $\eta(\mathfrak{s}_{a})=\mathfrak{s}_{a}$.
Note also that for any open subset $U$ of $X_{\mathscr{A}}$ the map
$\mathscr{A}\to{\mathcal{P}}_{\mathscr{A}}(U),\qquad a\mapsto\mathfrak{s}_{a}|_{U}$
is a surjective map of $C^{\infty}$-rings with the kernel $\bigcap_{x\in U}I_{x}$ (the ideals $I_{x}$ are given by (3.3)). Thus
${\mathcal{P}}_{\mathscr{A}}(U)\simeq\mathscr{A}/\bigcap_{x\in U}I_{x}.$
Finally note that for any $C^{\infty}$-ring $\mathscr{A}$ we get a map of $C^{\infty}$-rings
$\mathscr{A}\to{\mathcal{P}}(X_{\mathscr{A}}),\qquad a\mapsto\mathfrak{s}_{a}.$
which we may think of as a map
$\varepsilon_{\mathscr{A}}:\mathscr{A}\to\Gamma(\mathscr{O}_{\mathscr{A}}),\qquad\varepsilon_{\mathscr{A}}(a)=\mathfrak{s}_{a}.$
In general the map $\varepsilon_{\mathscr{A}}$ is neither injective nor surjective. For example, $\ker\varepsilon_{\mathscr{A}}=\bigcap_{x\in X_{\mathscr{A}}}I_{x}$, which may be nonzero.
Remark 3.16.
In the case where the $C^{\infty}$-ring $\mathscr{A}$ in Construction 3.14 is the ring $C^{\infty}(M)$ of smooth functions on a manifold $M$, the space $X_{\mathscr{A}}$ of points of $\mathscr{A}$ is canonically homeomorphic to the topological space $M$ underlying the manifold $M$. More generally, as a local $C^{\infty}$-ringed space $\operatorname{Spec}(C^{\infty}(M))$ is isomorphic to $(M,C^{\infty}_{M})$ where $C^{\infty}_{M}$ denotes the sheaf of smooth functions on the manifold $M$. Moreover there is a unique isomorphism $\underline{\varphi}_{M}:(M,C^{\infty}_{M})\to\operatorname{Spec}(C^{\infty}(M))$ that makes the diagram
${C^{\infty}(M)}$${\Gamma(C^{\infty}_{M})}$${\Gamma(\mathscr{O}_{C^{\infty}(M)})}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{\varepsilon_{C^{\infty}(M)}}$$\scriptstyle{\Gamma(\underline{\varphi}_{M})}$
commute. From now on we suppress the isomorphism $\underline{\varphi}_{M}:(M,C^{\infty}_{M})\to\operatorname{Spec}(C^{\infty}(M))$ for all manifolds $M$. That is, in effect, we set
$\operatorname{Spec}(C^{\infty}(M))=(M,C^{\infty}_{M})$
for all manifolds $M$.
Lemma 3.17.
Let $\mathscr{A}$ be a Poisson $C^{\infty}$-ring and $x:\mathscr{A}\to\mathbb{R}$ an $\mathbb{R}$-point. Then the ideal $I_{x}$ defined by (3.3) is a Poisson ideal.
Proof.
We need to show that for any $a\in I_{x}$ and any $b\in\mathscr{A}$ the bracket $\{a,b\}$ is again in $I_{x}$.
Since $a\in I_{x}$ there is an element $d\in\mathscr{A}$ so that $x(d)\not=0$ and $ad=0$ (cf. Remark 3.12. And then
$0=\{b,ad\}=\{b,a\}d+a\{b,d\}.$
Since $a\in I_{x}$ and $I_{x}$ is an ideal, $a\{b,d\}\in I_{x}$. Consequently $\{b,a\}d\in I_{x}$. Therefore there is $d^{\prime}\in\mathscr{A}$ with $x(d^{\prime})\not=0$ and $\{b,a\}dd^{\prime}=0$. Since $0\not=x(d)x(d^{\prime})=x(dd^{\prime})$, it follows that $\{b,a\}\in I_{x}$ and we are done. ∎
Theorem 3.18.
Let $\mathscr{A}$ be a Poisson $C^{\infty}$-ring (Definition 2.5). Then the structure sheaf $\mathscr{O}_{\mathscr{A}}$ of the affine $C^{\infty}$-scheme $\operatorname{Spec}(\mathscr{A})=(X_{\mathscr{A}},\mathscr{O}_{\mathscr{A}})$ is a sheaf of Poisson $C^{\infty}$-rings. That is, $\operatorname{Spec}(\mathscr{A})$ is a Poisson $C^{\infty}$-scheme.
Proof.
Since the structure sheaf $\mathscr{O}_{\mathscr{A}}$ is the sheafification of the presheaf ${\mathcal{P}}_{\mathscr{A}}$, it is enough to show that for every open set $U\subset X_{\mathscr{A}}$ the $C^{\infty}$-ring ${\mathcal{P}}_{\mathscr{A}}(U)$ is Poisson and that for pairs of open sets $V\subseteq U$ the restriction maps ${\mathcal{P}}_{\mathscr{A}}(U)\to{\mathcal{P}}_{\mathscr{A}}(V)$ are Poisson. Recall from Remark 3.15 that for every open $U\subset X_{\mathscr{A}}$ the map
$\mathscr{A}\to{\mathcal{P}}(U),\qquad a\mapsto\mathfrak{s}_{a}|_{U}$
is surjective with the kernel $\bigcap_{x\in U}I_{x}$. By Lemma 3.17 the ideals $I_{x}$ are Poisson for all $\mathbb{R}$-points $x$. Hence the intersection $\bigcap_{x\in U}I_{x}$ is a Poisson ideal, and therefore the $C^{\infty}$-ring ${\mathcal{P}}(U)$ is Poisson. Note that the bracket on ${\mathcal{P}}(U)$ is given by
$\{\mathfrak{s}_{a}|_{U},\mathfrak{s}_{b}|_{U}\}=\mathfrak{s}_{\{a,b\}}|_{U}.$
It follows that the restriction maps
${\mathcal{P}}_{\mathscr{A}}(U)\to{\mathcal{P}}_{\mathscr{A}}(V),\qquad\mathfrak{s}_{a}|_{U}\mapsto(\mathfrak{s}_{a}|_{U})|_{V}=\mathfrak{s}_{a}|_{V}$
are Poisson. ∎
Notation 3.19.
Poisson $C^{\infty}$-rings form a category, which we denote by $\mathsf{P}C^{\infty}\mathsf{Ring}$. The objects of $\mathsf{P}C^{\infty}\mathsf{Ring}$ are Poisson $C^{\infty}$-rings, the morphisms are maps of $C^{\infty}$-rings that preserve brackets.
Notation 3.20.
Poisson local $C^{\infty}$-ringed spaces also from a category: the objects are Poisson local $C^{\infty}$-ringed spaces. We denote it by $\mathsf{PLC^{\infty}RS}$. A map in this category from a Poisson local $C^{\infty}$-ringed space $(X,\mathscr{O}_{X})$ to another Poisson local $C^{\infty}$-ringed space $(Y,\mathscr{O}_{Y})$ is a map $\underline{f}=(f,f_{\#}):(X,\mathscr{O}_{X})\to(Y,\mathscr{O}_{Y})$ of local $C^{\infty}$-ringed spaces so that $f_{\#}:\mathscr{O}_{Y}\to f_{*}\mathscr{O}_{X}$ is a map of sheaves of Poisson $C^{\infty}$-rings. That is, for every open set $U\subset X$ the map
$f_{\#,U}:\mathscr{O}_{Y}(U)\to\mathscr{O}_{X}(f^{-1}(U))$
of $C^{\infty}$-rings preserves the Poisson bracket.
We can restrict Dubuc’s spectrum functor $\operatorname{Spec}:{C^{\infty}\mathsf{Ring}}^{\mbox{\sf{\tiny{op}}}}\to\mathsf{LC^{\infty}RS}$ to the subcategory of $\mathsf{P}C^{\infty}\mathsf{Ring}$ of Poisson $C^{\infty}$-rings. We expect the image of $\operatorname{Spec}|_{\mathsf{P}C^{\infty}\mathsf{Ring}}$ to land in the category $\mathsf{PLC^{\infty}RS}$ of Poisson local $C^{\infty}$-ringed spaces. To prove this we need to recall Joyce’s construction of $\operatorname{Spec}$ on morphisms. As a first step we recall a lemma, which is well-known to experts.
Lemma 3.21.
Let $\varphi:\mathscr{A}\to\mathscr{B}$ be a map of $C^{\infty}$-rings and $x:\mathscr{B}\to\mathbb{R}$ a point. Then $\varphi$ induces a unique map $\varphi_{x}:\mathscr{A}_{x\circ\varphi}\to\mathscr{B}_{x}$ of $C^{\infty}$-rings on their respective localizations making the diagram
(3.7) ${\mathscr{A}}$${\mathscr{B}}$${\mathscr{A}_{x\circ\varphi}}$${\mathscr{B}_{x}}$$\scriptstyle{\pi_{x\circ\varphi}}$$\scriptstyle{\pi_{x}}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi_{x}}$
commute.
Proof.
Since $\pi_{y}:\mathscr{A}\to\mathscr{A}_{y}$ is a localization of the $C^{\infty}$-ring $\mathscr{A}$ at the set $\{y\not=0\}:=\{a\in\mathscr{A}\mid y(a)\not=0\}$ and $\pi_{x}:\mathscr{B}\to\mathscr{B}_{x}$ is the localization of $\mathscr{B}$ at $\{x\not=0\}$, it is enough to show that
(3.8) $\varphi(\{x\circ\varphi\not=0\})\subseteq\{x\not=0\}.$
Since
$(x\circ\varphi)(a)\not=0\quad\Leftrightarrow\quad x(\varphi(a))\not=0,$
(3.8) holds and we are done. ∎
Construction 3.22 (Construction of $\operatorname{Spec}$ on morphisms).
Let $\varphi:\mathscr{A}\to\mathscr{B}$ be a map of $C^{\infty}$-rings. We construct $\operatorname{Spec}(\varphi):\operatorname{Spec}(\mathscr{B})\to\operatorname{Spec}(\mathscr{A})$, a map of local $C^{\infty}$-ringed spaces. First define a map of sets $f(\varphi):X_{\mathscr{B}}\to X_{\mathscr{A}}$ (where as before $X_{\mathscr{A}}=\operatorname{Hom}(\mathscr{A},\mathbb{R})$ and $X_{\mathscr{B}}=\operatorname{Hom}(\mathscr{B},\mathbb{R})$) by
$f(\varphi)(x):=x\circ\varphi.$
Recall that $\{V_{a}=\{y\in X_{\mathscr{A}}\mid y(a)\not=0\}\}_{a\in\mathscr{A}}$ is a basis for the topology on $X_{\mathscr{A}}$ and similarly $\{U_{b}=\{x\in X_{\mathscr{B}}\mid x(b)\not=0\}\}_{b\in\mathscr{B}}$ is a basis for the topology on $X_{\mathscr{B}}$. Since
$(f(\varphi))^{-1}(V_{a})=\{x\in X_{\mathscr{B}}\mid f(\varphi)x\in V_{a}\}=\{x\in X_{\mathscr{B}}\mid x(\varphi(a))\not=0\}=U_{\varphi(a)},$
the map $f(\varphi)$ is continuous with respect to the Zariski topologies on $X_{\mathscr{B}}$ and $X_{\mathscr{A}}$. It remains to construct a map of sheaves
$f(\varphi)_{\#}:\mathscr{O}_{\mathscr{A}}\to f(\varphi)_{*}\mathscr{O}_{\mathscr{B}}.$
Fix an open set $U\subseteq X_{\mathscr{A}}$. We construct a map
$f(\varphi)_{\#,U}:{\mathcal{P}}_{\mathscr{A}}(U)\to\mathscr{O}_{\mathscr{B}}(f(\varphi)^{-1}U).$
of $C^{\infty}$-rings as follows. Recall that
${\mathcal{P}}_{\mathscr{A}}(U)=\{\mathfrak{s}_{a}:U\to\bigsqcup_{x\in U}\mathscr{A}_{x}\mid a\in\mathscr{A}\}$
where, as in (3.4), $\mathfrak{s}_{a}(y)=a_{y}$ for all $y\in U$. Here and below to reduce the clutter we write $\mathfrak{s}_{a}$ when we mean $\mathfrak{s}_{a}|_{U}$. Given $\mathfrak{s}_{a}\in{\mathcal{P}}_{\mathscr{A}}(U)$ consider
$\tilde{\mathfrak{s}}_{a}:f(\varphi)^{-1}(U)\to\bigsqcup_{x\in f(\varphi)^{-1}(U)}\mathscr{B}_{x},\qquad\tilde{\mathfrak{s}}_{a}(x):=\varphi_{x}(\mathfrak{s}_{a}(f(\varphi)x)),$
where $\varphi_{x}:\mathscr{A}_{f(\varphi)x}\to\mathscr{B}_{x}$ is the map from Lemma 3.21. Note that
$\varphi_{x}(\mathfrak{s}_{a}(f(\varphi)x))=\varphi_{x}(\pi_{f(\varphi)x}a)=\pi_{x}(\varphi(a)),$
where the last equality is commutativity of (3.7). Hence $\tilde{\mathfrak{s}}_{a}=\mathfrak{s}_{\varphi(a)}\in{\mathcal{P}}_{\mathscr{B}}(f(\varphi)^{-1}(U))\subset\mathscr{O}_{\mathscr{B}}(f(\varphi)^{-1}(U))$. We therefore define
$f(\varphi)_{\#,U}:{\mathcal{P}}_{\mathscr{A}}(U)\to\mathscr{O}_{\mathscr{B}}(f(\varphi)^{-1}U)$
by
$f(\varphi)_{\#,U}(\mathfrak{s}_{a}):=\mathfrak{s}_{\varphi(a)}$
for all $a\in\mathscr{A}$. It is easy to see that the family of maps $\{f(\varphi)_{\#,U}\}_{U\subset X_{\mathscr{A}}}$ is a map of presheaves $f(\varphi)_{\#}:{\mathcal{P}}_{\mathscr{A}}\to f(\varphi)_{*}\mathscr{O}_{\mathscr{B}}$. By the universal property of sheafification we get a map of sheaves $f(\varphi)_{\#}:\mathscr{O}_{\mathscr{A}}\to f(\varphi)_{*}\mathscr{O}_{\mathscr{B}}$.
We define
$\operatorname{Spec}(\varphi)\equiv\underline{f(\varphi)}:=(f(\varphi),f(\varphi)_{\#}).$
Lemma 3.23.
For any map $\varphi:\mathscr{A}\to\mathscr{B}$ of Poisson $C^{\infty}$-rings the induced map on spectra $\operatorname{Spec}(\varphi)\equiv(f(\varphi),f(\varphi)_{\#}):\operatorname{Spec}(\mathscr{B}))\to\operatorname{Spec}(\mathscr{A})$ is Poisson.
Proof.
It is enough to show that for any open subset $U$ of $X_{\mathscr{B}}$ the map
$f(\varphi)_{\#,U}:{\mathcal{P}}_{\mathscr{A}}(U)\to\mathscr{O}_{\mathscr{B}}(f(\varphi)^{-1}U),\qquad\mathfrak{s}_{a}\mapsto\mathfrak{s}_{\varphi(a)}$
is Poisson. Recall that the bracket on ${\mathcal{P}}_{\mathscr{A}}(U)$ is given by
$\{\mathfrak{s}_{a},\mathfrak{s}_{b}\}=\mathfrak{s}_{\{a,b\}}$
for all $a,b\in\mathscr{A}$ (as before we drop the restrictions to $U$ to reduce the clutter). Since $\varphi(\{a,b\})=\{\varphi(a),\varphi(b)\}$ the result follows. ∎
4. Applications
4.i. Symplectic reduction
Let $(M,\omega,\mu:M\to\mathfrak{g}^{\vee})$ be a symplectic manifold with a Hamiltonian action of a Lie group $G$ and associated equivariant moment map $\mu$. Let $\mathscr{A}$ be the $C^{\infty}$-ring of $G$-invariant functions $C^{\infty}(M)^{G}$. As we observed in Example 2.17, for any point $\alpha\in\mathfrak{g}^{\vee}$ the ideal $I_{\alpha}\subset\mathscr{A}$ of invariant functions that vanish on the level set $\mu^{-1}(\alpha)$ is a Poisson ideal. Hence $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ is a Poisson $C^{\infty}$-scheme.
In the case where the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is sufficiently nice and the action of $G$ on $M$ is proper the space $X_{\mathscr{A}/I_{\alpha}}$ of points of the scheme $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ is homeomorphic to $\mu^{-1}(\alpha)/G_{\alpha}$ (where $G_{\alpha}$ denotes the stabilizer of $\alpha\in\mathfrak{g}^{\vee}$ under the coadjoint actions). We’ll discuss it in the next section, Section A. Therefore we may regard the scheme $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ as (a possibly singular) symplectic quotient (a.k.a. reduced space) $M/\!/_{\alpha}G$.
If the action of $G$ on $M$ is not proper, the space $X_{\mathscr{A}/I_{\alpha}}$ may be quite different from the topological quotient $\mu^{-1}(\alpha)/G_{\alpha}$. Here is an example.
Example 4.1.
Consider the action of $\mathbb{R}$ on the torus $\mathbb{T}^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$ by an irrational flow:
$t\cdot[(x,y)]:=[x+t,y+\sqrt{2}t]$
for all $t\in\mathbb{R}$ and all $[x,y]\in\mathbb{R}^{2}/\mathbb{Z}^{2}$. Then the lifted action of $\mathbb{R}$ on the cotangent bundle $T^{*}\mathbb{T}^{2}\simeq\mathbb{T}^{2}\times\mathbb{R}^{2}$ is Hamiltonian with the moment $\mu([x,y],p_{x},p_{y})=p_{x}+\sqrt{2}p_{y}$. Consequently
$\mu^{-1}(\alpha)=\{([x,y],p_{x},p_{y})\in T^{*}\mathbb{T}^{2}\mid p_{x}=\alpha-\sqrt{2}p_{y}\}$
and
$\mu^{-1}(\alpha)/\mathbb{R}\simeq(\mathbb{T}^{2}/\mathbb{R})\times\mathbb{R}.$
On the other hand the Poisson algebra of $\mathbb{R}$-invariant functions on $T^{*}\mathbb{T}^{2}$ is $C^{\infty}(T^{*}\mathbb{T}^{2})^{\mathbb{R}}\simeq C^{\infty}(\mathbb{R}^{2})$ with the zero Poisson bracket and
$I_{\alpha}\simeq\{f\in C^{\infty}(\mathbb{R}^{2})\mid f(\alpha-\sqrt{2}p_{y},p_{y})=0\textrm{ for all }p_{y}\in\mathbb{R}\}.$
Hence
$C^{\infty}(T^{*}\mathbb{T}^{2})^{\mathbb{R}}\simeq C^{\infty}(\mathbb{R}^{2})/I_{\alpha}\simeq C^{\infty}(\mathbb{R})$
(with the zero Poisson bracket) and
$X_{C^{\infty}(T^{*}\mathbb{T}^{2})^{\mathbb{R}}}\simeq X_{C^{\infty}(\mathbb{R}^{2})/I_{\alpha}}\simeq X_{C^{\infty}(\mathbb{R})}=\mathbb{R}.$
Note that there is an evident map
$\mu^{-1}(\alpha)/\mathbb{R}\simeq(\mathbb{T}^{2}/\mathbb{R})\times\mathbb{R}\to X_{C^{\infty}(T^{*}\mathbb{T}^{2})^{\mathbb{R}}}\simeq\mathbb{R}$
which identifies all the points that cannot be separated by invariant functions.
4.ii. Poisson reduction
The construction of the Poisson scheme in 4.i does not require the assumption that the manifold $M$ is symplectic — we only need that $M$ is a Poisson manifold, that the action of the group $G$ preserves the Poisson bracket (and hence the $C^{\infty}$-subgring $C^{\infty}(M)^{G}$ of invariant functions is Poisson) and that the Hamiltonian flows of invariant functions preserve a closed subset $X\subset M$. Then by Lemma 2.14 the vanishing ideal
${{\mathcal{I}}}(X):=\{f\in C^{\infty}(M)^{G}:f|_{X}=0\}$
is a Poisson ideal in the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ of invariant functions. Hence $C^{\infty}(M)^{G}/{\mathcal{I}}(X)$ is a Poisson $C^{\infty}$-ring and $\operatorname{Spec}(C^{\infty}(M)^{G}/{\mathcal{I}}(X)$ is a Poisson $C^{\infty}$-scheme.
For example, the action of the group $G$ on the Poisson manifold $M$ may have an equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$ [Et]. Then as in the symplectic case the level sets $\mu^{-1}(\alpha)$, $\alpha\in\mathfrak{g}^{\vee}$ are preserved by the Hamiltonian flows of invariant functions and we may take the closed set $X$ to be a level set $\mu^{-1}(\alpha)$, just as we did in the symplectic case.
Since $\varpi:C^{\infty}(M)^{G}\to C^{\infty}(M)^{G}/{\mathcal{I}}(X)$ is surjective, Lemma A.5 implies that the induced map $f(\varphi):X_{C^{\infty}(M)^{G}/{\mathcal{I}}(X)}\to X_{C^{\infty}(M)^{G}}$ of spaces is a closed embedding. Furthermore, by Lemma A.10, if the action of $G$ on $M$ is proper then the space of points $X_{C^{\infty}(M)^{G}}$ is homeomorphic to $M/G$. Hence in the case of proper actions the space of points of the affine Poisson scheme $\operatorname{Spec}(C^{\infty}(M)^{G}/{\mathcal{I}}(X))$ is homeomorphic to a closed subset of the orbit space $M/G$.
4.iii. quasi-Poisson reduction
Let $(M,\mathsf{P},\Phi:M\to G)$ be a quasi-Poisson manifold in the sense of Alekseev, Kosmann-Scwarzbach and Meinrenken [AKSM]. As explained in Section 6 of [AKSM] the $C^{\infty}$-ring of invariant functions $C^{\infty}(M)^{G}$ is a Poisson $C^{\infty}$-ring. Therefore $\operatorname{Spec}(C^{\infty}(M)^{G})$ is a Poisson $C^{\infty}$-scheme.
It is shown in [AKSM] in the course of the proof of Theorem 6.1 that for any invariant function $f$ on $M$ the moment map $\Phi$ is constant along the integral curves of the Hamiltonian vector field of $f$. Arguing as in the proof of Lemma 2.14 we see that for any conjugacy class $\mathscr{C}\subset G$ the ideal
$I_{\mathscr{C}}:=\{f\in C^{\infty}(M)^{G}\mid f|_{\Phi^{-1}(\mathscr{C})}=0\}$
is a Poisson ideal in the Poisson $C^{\infty}$-ring $C^{\infty}(M)^{G}$. Consequenty the quotient $C^{\infty}(M)^{G}/I_{\mathscr{C}}$ is a Poisson $C^{\infty}$-ring and its spectrum $\operatorname{Spec}(C^{\infty}(M)^{G}/I_{\mathscr{C}})$ is a Poisson $C^{\infty}$-scheme. As before the quotient map $C^{\infty}(M)^{G}\to C^{\infty}(M)^{G}/I_{\mathscr{C}}$ induces a closed embedding $X_{C^{\infty}(M)^{G}/I_{\mathscr{C}}}\hookrightarrow X_{C^{\infty}(M)^{G}}$.
4.iv. Schemes associated to various Ideals in Poisson manifolds
Let $P$ be a Poisson manifold and $I\subset C^{\infty}(P)$ an ideal in the $C^{\infty}$-ring $C^{\infty}(P)$. By Lemma 2.20 the normalizer of $I$ in the Poisson algebra $C^{\infty}(P)$ is a Poisson $C^{\infty}$-ring. We now consider two cases.
Case a. The ideal $I$ is a Poisson subalgebra of $C^{\infty}(P)$, i.e, for all $f,g\in I$ the bracket $\{f,g\}\in I$. Then the ideal $I$ is contained in its normalizer $N(I)$. Hence the quotient $N(I)/I$ is a Poisson $C^{\infty}$-ring and its spectrum $\operatorname{Spec}(N(I)/I$ is a Poisson $C^{\infty}$-scheme. This is a version of coisotropic reduction in Poisson geometry.
In more detail, the inclusion $N(I)\hookrightarrow C^{\infty}(P)$ induces an injective map $\varphi:N(I)/I\to C^{\infty}(P)/I$. Taking $\operatorname{Spec}$ we get a map of $C^{\infty}$-schemes $\operatorname{Spec}(\varphi):\operatorname{Spec}(C^{\infty}(P)/I)\to\operatorname{Spec}(N(I)/I)$.
Note that in general if $\psi:\mathscr{A}\to\mathscr{B}$ is an injective map of $C^{\infty}$-rings, the induced map of points $f(\psi):X_{\mathscr{B}}\to X_{\mathscr{A}}$ need not be surjective. For example consider a smooth map $F:M\to N$ between two manifolds with $F(M)$ dense in $N$. Then $F^{*}:C^{\infty}(N)\to C^{\infty}(M)$ is injective and $f(F^{*})=F$ (we use the notation of Construction 3.22), so $f(F^{*})$ it’s not surjective.
However, suppose $C\subset P$ is a coisotropic submanifold and let $I$ denote the ideal of functions that vanish on $C$. Then $I$ is a Poisson subalgebra of $C^{\infty}(P)$ and the Poisson $C^{\infty}$-ring $N(I)/I$ is isomorphic to the $C^{\infty}$-ring of functions on $C$ that are constant along the leaves of the null foliation $\mathscr{F}$. If $C/\mathscr{F}$ is a manifold then it is isomorphic to $\operatorname{Spec}(N(I)/I)$. In general we view the Poisson $C^{\infty}$-scheme $\operatorname{Spec}(N(I)/I)$ as the coisotropic reduction of $C$. Note that the underlying topological space $X_{N(I)/I}$ may be nicer than the leaf space $C/\mathscr{F}$ since (as was mentioned above) the space of points of a $C^{\infty}$-ring is Hausdorff and completely regular. We have seen an instance of it when we looked at the reduction of the cotangent bundle of the 2-torus $T^{*}\mathbb{T}$ by the irrational flow (Example 4.1).
Example 4.2 (Sniatycki-Weinstein reduction).
We now revisit Example 2.21. Recall that one starts with a symplectic manifold $(M,\omega)$ equipped with a Hamiltonian action of a connected Lie group $G$ and associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Then the ideal $\mathscr{I}\subset C^{\infty}(M)$ generated by the components of the moment map $\mu$ is closed under the Poisson brackets, hence $N(\mathscr{I})/\mathscr{I}$ is Poisson $C^{\infty}$-ring equipped with an injective map $\varphi:N(\mathscr{I})/\mathscr{I}\to C^{\infty}(M)/\mathscr{I}$. Hence $\operatorname{Spec}(N(\mathscr{I})/\mathscr{I}$ is a Poisson scheme which we may view as a coisotropic reduction of $C^{\infty}(M)/\mathscr{I}$.
Note that since the quotient map $q:C^{\infty}(M)\to C^{\infty}(M)/\mathscr{I}$ is surjective, the induced map $f(q):X_{C^{\infty}(M)/\mathscr{I}}\to X_{C^{\infty}(M)}\simeq M$ is a closed embedding by Lemma A.5. However, in general the ideal $\mathscr{I}$ may be strictly smaller than the vanishing ideal of $\mu^{-1}(0)$ (cf. [AGJ]) and the scheme $\operatorname{Spec}(C^{\infty}(M)/\mathscr{I})$ need not be reduced.
Case b. The ideal $I$ is an arbitrary ideal of the $C^{\infty}$-ring $C^{\infty}(P)$. Then the intersection $N(I)\cap I\subseteq N(I)$ is a Poisson ideal in $N(I)$. Consequently $\operatorname{Spec}(N(I)/I\cap N(I))$ is a Poisson scheme. Note that the inclusion $N(I)\hookrightarrow C^{\infty}(P)$ induces an injective map
(4.1) $\varphi:N(I)/I\cap N(I)\hookrightarrow C^{\infty}(P)/I.$
If the map $\varphi$ above is an isomorphism then since $N(I)/I\cap N(I)$ is Poisson $C^{\infty}$-ring, the quotient $C^{\infty}(P)/I$ is a also Poisson $C^{\infty}$-ring. Then the close subscheme $\operatorname{Spec}(C^{\infty}(P)/I)$ of the manifold $P$ is a Poisson scheme. When $I$ is a vanishing ideal of a submanifold $M\subset P$, then $\operatorname{Spec}(C^{\infty}(P)/I)=M$ and $M$ is a Poisson-Dirac submanifold of $P$.
In general whenever $\varphi$ is an isomorphism we can define $\operatorname{Spec}(C^{\infty}(P)/I)$ to be a Poisson-Dirac subscheme $P$ (without any smoothness assumption).
If the map $\varphi$ fails to be surjective we still have a map
$\operatorname{Spec}(\varphi):\operatorname{Spec}(C^{\infty}(P)/I)\to\operatorname{Spec}N(I)/I\cap N(I)$
of local $C^{\infty}$-ringed spaces. We can view the map as a kind of coisotropic reduction of the closed subscheme $\operatorname{Spec}(C^{\infty}(P)/I)$ of the Poisson manifold $P$.
Appendix A Computing $\operatorname{Spec}(\mathscr{A})$
We start by recalling what it means for a $C^{\infty}$-ring $\mathscr{A}$ to be finitely generated. For such rings Dubuc [Du] gave a fairly concrete description of $\operatorname{Spec}(\mathscr{A})$ which we describe. Next we consider point determined $C^{\infty}$-rings (see Definition A.2). We show that if $\mathscr{A}$ is point determined then the structure sheaf $\mathscr{O}_{\mathscr{A}}$ of $\operatorname{Spec}(\mathscr{A})$ is a sheaf of functions on the space $X_{\mathscr{A}}$ of its points. For finitely generated and point determined $C^{\infty}$-rings the description of $\operatorname{Spec}(\mathscr{A})$ simplifies considerably: $\operatorname{Spec}(\mathscr{A})$ is a closed subset of some $\mathbb{R}^{N}$ with the induced sheaf of “smooth” functions.
We then study $\operatorname{Spec}(\mathscr{A})$ where $\mathscr{A}$ is a ring of invariant functions $C^{\infty}(M)^{G}$ for an action of a Lie group $G$ on a manifold $M$. Such rings are point determined and there is always a map $\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M)^{G}}$ from the orbit space to the space of points of $C^{\infty}(M)^{G}$. If the invariant functions separate points the map $\overline{\operatorname{\mathsf{ev}}}$ is injective. If $C^{\infty}(M)^{G}$ contains a proper function then the map $\overline{\operatorname{\mathsf{ev}}}$ is surjective. If the action of $G$ on $M$ is proper then $\overline{\operatorname{\mathsf{ev}}}$ is a homeomorphism. And then the structure sheaf of $\operatorname{Spec}(C^{\infty}(M)^{G})$ is the sheaf of functions on the orbit space $M/G$ which has a familiar description.
Finally we consider the case where the action of $G$ on $M$ is proper and the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is finitely generated. This happens, for example, when $G$ is compact and the action has finitely many orbit type by a deep theorem of Schwarz [Sch]. Note that finiteness of the number of orbit types is not necessary — see Example 6.9 in [KL1]. We then have two descriptions of $\operatorname{Spec}(C^{\infty}(M)^{G}$: as a sheaf on functions on $M/G$ and as sheaf of functions on a closed subset $Z$ of some $\mathbb{R}^{n}$. Not surprisingly the two descriptions a related in a way that should look familiar. We now get to work.
A.i. $\operatorname{Spec}(\mathscr{A})$ for a finitely generated $C^{\infty}$-ring $\mathscr{A}$
Recall that for any natural number $N$ the $C^{\infty}$-ring $C^{\infty}(\mathbb{R}^{N})$ is freely generated by the set $\{x_{1},\ldots,x_{N}\}$ of the standard coordinate functions (see [MR], for example). This is because for any smooth function $f\in C^{\infty}(\mathbb{R}^{N})$
$f=f\circ(x_{1},\ldots,x_{N})=f_{C^{\infty}(\mathbb{R}^{N})}(x_{1},\ldots,x_{N}).$
Definition A.1.
A $C^{\infty}$-ring $\mathscr{A}$ is finitely generated if there is a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to\mathscr{A}$ of $C^{\infty}$-rings. (And then $\varpi(x_{1}),\ldots\varpi(x_{N})\in\mathscr{A}$ are the generators of $\mathscr{A}$.)
In [Du] on p. 687 in the paragraph right below Proposition 12 Dubuc describes the affine scheme $\operatorname{Spec}(\mathscr{A})$ for a finitely generated $C^{\infty}$-ring $\mathscr{A}$ as follows. Since $\mathscr{A}$ is finitely generated there is a surjective map of $C^{\infty}$-rings $\Pi:C^{\infty}(\mathbb{R}^{n})\to\mathscr{A}$. Let $J$ denote the kernel of $\Pi$ and let
(A.1) $Z=Z_{J}:=\{p\in\mathbb{R}^{n}\mid f(p)=0\textrm{ for all }f\in J\}$
denote the zero set of the ideal $J$. For each open set $U\subset\mathbb{R}^{n}$ we then have the quotient $C^{\infty}$-ring
${\mathcal{P}}(U)=C^{\infty}(U)/J_{U}$
where $J_{U}\subset C^{\infty}(U)$ is the ideal generated by the set $J|_{U}=\{f|_{U}\mid f\in J\}$. Note that if $U\cap Z_{J}=\varnothing$, then ${\mathcal{P}}(U)=0$. Let ${\mathcal{E}}$ be the sheafification of ${\mathcal{P}}$. The support of ${\mathcal{E}}$ is $Z_{J}$, hence we can view ${\mathcal{E}}$ as a sheaf on $Z$. Dubuc then tells us that $\operatorname{Spec}(\mathscr{A})\simeq(Z,{\mathcal{E}})$.
A.ii. $\operatorname{Spec}(\mathscr{A})$ for a point determined $C^{\infty}$-ring $\mathscr{A}$
It will be useful for us to formulate the definition of a point determined $C^{\infty}$-ring as follows (it is equivalent to Definition 4.1(a) in [MR, p. 44]).
Definition A.2.
A $C^{\infty}$-ring $\mathscr{A}$ is point determined if for every element $a\in\mathscr{A}$
$x(a)=0\textrm{ for all }x\in\operatorname{Hom}(\mathscr{A},\mathbb{R})\equiv X_{\mathscr{A}}\quad\Rightarrow a=0.$
Equivalently the map
$\mathscr{A}\to\prod_{x\in X_{\mathscr{A}}}\mathbb{R},\quad a\mapsto(x(a))_{x\in X_{\mathscr{A}}}$
is an embedding.
Here is another interpretation of what it means for a $C^{\infty}$-ring $\mathscr{A}$ to be point determined which is more geometrically intuitive. For every element $a\in\mathscr{A}$ we have a function $a_{*}:X_{\mathscr{A}}\to\mathbb{R}$, $a_{*}(x):=x(a)$. The functions $a_{*}$ are continuous with respect to the Zariski topology $\mathscr{T}_{\mathscr{A}}$ on $X_{\mathscr{A}}$ and, moreover, $\mathscr{T}_{\mathscr{A}}$ is the smallest topology making all the funtions $a_{*}$ continuous [J]. In general the map
(A.2) $\mathscr{A}\to C^{0}(X_{\mathscr{A}}),\quad a\mapsto a_{*}$
may not be injective. It is injective if and only if $\mathscr{A}$ is point determined.
We now argue that if a $C^{\infty}$-ring $\mathscr{A}$ is point determined then the structure sheaf $\mathscr{O}_{\mathscr{A}}$ of the affine $C^{\infty}$-scheme $\operatorname{Spec}(\mathscr{A})$ is a sheaf of continuous functions on its space $X_{\mathscr{A}}$ of points.
Lemma A.3.
Let $\mathscr{A}$ be a point determined $C^{\infty}$-ring. Define a presheaf ${\mathcal{P}}$ on the space $X_{\mathscr{A}}$ of points of $\mathscr{A}$ by
(A.3) ${\mathcal{P}}(U)=\{a_{*}|_{U}\mid a\in\mathscr{A}\}$
for all $U\in\mathsf{Open}(X_{\mathscr{A}})$ (where as before $a_{*}(x):=x(a)$ for all $a\in\mathscr{A}$ and all $x\in X_{\mathscr{A}}$) . The structure sheaf $\mathscr{O}_{\mathscr{A}}$ of $\operatorname{Spec}(\mathscr{A})$ is isomorphic to the sheafification of the presheaf ${\mathcal{P}}$.
Proof.
Recall that the structure sheaf $\mathscr{O}_{\mathscr{A}}$ is the sheafification of the presheaf ${\mathcal{P}}_{\mathscr{A}}$ given by
${\mathcal{P}}_{\mathscr{A}}(U)=\{\mathfrak{s}_{a}|_{U}\mid a\in\mathscr{A}\}$
(cf. (3.6)) and that ${\mathcal{P}}_{\mathscr{A}}(U)\simeq\mathscr{A}/\bigcap_{x\in U}I_{x}$. We construct a map of presheaves $\psi:{\mathcal{P}}_{\mathscr{A}}\to{\mathcal{P}}$ and prove that $\psi$ is an isomorphism.
Given $U\in\mathsf{Open}(X_{\mathscr{A}})$ consider the map
$\mu_{U}:\mathscr{A}\to{\mathcal{P}}(U),\quad\mu_{U}(a):=a_{*}|_{U}.$
Suppose $a\in\bigcap_{x\in U}I_{x}$. Then for any $x\in U$, we have $a\in I_{x}$, which means that there is $d\in\mathscr{A}$ so that $x(d)\not=0$ and $ad=0$. Since $0\not=x(d)=d_{*}(x)$ and $d_{*}$ is continuous there is a neighborhood $V$ of $x$ so that $d_{*}(y)\not=0$ for all $y\in V$. Since $ad=0$, $0=(ad)_{*}=a_{*}d_{*}$. Hence $a_{*}|_{V}=0$. Since $x\in U$ is arbitrary, $a_{*}|_{U}=0$. Therefore
$\psi_{U}:{\mathcal{P}}_{\mathscr{A}}(U)\simeq\mathscr{A}/\bigcap_{x\in U}I_{x}\to{\mathcal{P}}(U),\quad\mathfrak{s}_{a}|_{U}\mapsto a_{*}|_{U}$
is a well-defined map of $C^{\infty}$-rings. It is not hard to check that the collection of maps $\{\psi_{U}\}_{U\in\mathsf{Open}(X_{\mathscr{A}})}$ assemble into a map of presheaves $\psi:{\mathcal{P}}_{\mathscr{A}}\to{\mathcal{P}}$.
It remains to show that the maps $\psi_{U}$ are isomorphism. It’s clear from the construction that $\psi_{U}:{\mathcal{P}}_{\mathscr{A}}(U)\to{\mathcal{P}}(U)$ is onto for every open set $U$. Suppose $\mathfrak{s}_{a}|_{U}\in\ker(\psi_{U})$. Then $a_{*}|_{U}=0$. Since the topology on $X_{\mathscr{A}}$ is the smallest topology such that all functions $a_{*}$ ($a\in\mathscr{A})$) are continuous, for every open set $V\subset X_{\mathscr{A}}$ for every $x\in V$ there is $d\in\mathscr{A}$ so that $d_{*}(x)\not=0$ and $\operatorname{supp}(d_{*})\subset V$ (see, for example [KL2, Lemma 2.15]). Then
$0=d_{*}a_{*}=(da)_{*}$
Since the map (A.2) is injective ($\mathscr{A}$ is point determned), it follows that $da=0$ Since $x(d)=x_{*}(d)\not=0$ it follows that $a\in I_{x}$. Since $x\in U$ is arbitrary, $a\in\bigcap_{x\in U}I_{x}$. Therefore $\mathfrak{s}_{a}|_{U}=0$ and $\psi_{U}$ is injective. ∎
We now describe $\operatorname{Spec}(\mathscr{A})$ when the $C^{\infty}$-ring $\mathscr{A}$ is finitely generated and point determined.
Lemma A.4.
Let $J\subset C^{\infty}(\mathbb{R}^{n})$ be an ideal and
$Z=Z_{J}=\{p\in C^{\infty}(\mathbb{R}^{n})\mid f(p)=0\textrm{ for all }f\in J\}$
is its vanishing set and
$I(Z_{J})=\{f\in C^{\infty}(\mathbb{R}^{n})\mid f(p)=0\textrm{ for all }p\in Z_{J}$
the zero ideal of $Z_{J}$.
If the quotient $\mathscr{A}:=C^{\infty}(\mathbb{R}^{n})/J$ is a point determined $C^{\infty}$-ring then
$I(Z_{J})=J.$
Proof.
Since $J\subset I(Z_{J})$ we only need to show that $I(Z_{J})\subset J$.
Recall that for every point $p\in\mathbb{R}^{n}$ we have an $\mathbb{R}$-point $\operatorname{\mathsf{ev}}_{p}:C^{\infty}(\mathbb{R}^{n})\to\mathbb{R}$ and that the map $\operatorname{\mathsf{ev}}:\mathbb{R}^{n}\to\operatorname{Hom}(C^{\infty}(\mathbb{R}^{n}),\mathbb{R})$ is a bijection (cf. Remark 3.4). The maps $\operatorname{\mathsf{ev}}_{p}$ factor through $C^{\infty}(\mathbb{R}^{n})/J$ if and only if $p\in Z_{J}$. We thus have an induced map
(A.4) $\overline{\operatorname{\mathsf{ev}}}:Z_{J}\to\operatorname{Hom}(C^{\infty}(\mathbb{R}^{n}/J),\mathbb{R}),\quad\overline{\operatorname{\mathsf{ev}}}_{p}(f+J)=f(p).$
If $x:C^{\infty}(\mathbb{R}^{n})/J\to\mathbb{R}$ is an $\mathbb{R}$-point then $x\circ\Pi:C^{\infty}(\mathbb{R}^{n})\to\mathbb{R}$ is also an $\mathbb{R}$-point (where $\Pi:C^{\infty}(\mathbb{R}^{n})\to C^{\infty}(\mathbb{R}^{n})/J$ is the quotient map). Consequently there is a unique $p\in\mathbb{R}^{n}$ so that
$x\circ\Pi=\operatorname{\mathsf{ev}}_{p}.$
Since $(x\circ\Pi)\,(f)=0$ for all $f\in J$, such point $p$ must lie in $Z_{J}$. Therefore $\overline{\operatorname{\mathsf{ev}}}:Z_{J}\to\operatorname{Hom}(C^{\infty}(\mathbb{R}^{n}/J),\mathbb{R})$ is surjective, i.e.,
$\operatorname{Hom}(\mathscr{A},\mathbb{R})=\{\overline{ev}_{p}\mid p\in Z_{J}\}.$
Suppose now $f\in I(Z_{J})$. Then
$0=f(p)=\operatorname{\mathsf{ev}}_{p}(f)=\overline{\operatorname{\mathsf{ev}}}_{p}(f+J)$
for all $p\in Z_{J}$. Hence $x(f+J)=0$ for all $x\in\operatorname{Hom}(\mathscr{A},\mathbb{R})$. Since $\mathscr{A}$ is point determined $f+J=0$ in $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})/J$. Thus $f\in J$. ∎
It follows from A.i that if $\mathscr{A}=C^{\infty}(\mathbb{R}^{n})/J$ is point determined then $X_{\mathscr{A}}=Z_{J}$ and $\mathscr{O}_{A}$ is given by
$\mathscr{O}_{\mathscr{A}}(U)=C^{\infty}(\widetilde{U})|_{U}$
for $U\in\mathsf{Open}(Z_{J})$. Here $\widetilde{U}\subset\mathbb{R}^{n}$ is any open set with $\widetilde{U}\cap Z_{J}=U$.
The fact that the space of points $X_{C^{\infty}(\mathbb{R}^{n})/J}$ embeds into $\mathbb{R}^{n}$ as a closed subspace is not a coincidence.
Lemma A.5.
Let $\varpi:\mathscr{B}\to\mathscr{A}$ be a surjective map of $C^{\infty}$-rings. Then the induced map on spaces $f(\varpi):X_{\mathscr{A}}\to X_{\mathscr{B}}$, $f(\varphi)x:=x\circ\varpi$ is a closed embedding.
Proof.
For an ideal $I\subset\mathscr{B}$ define the set of its zeros to be
$Z_{I}:=\{y\in X_{\mathscr{B}}\mid y(f)=0\textrm{ for all }f\in I\}.$
Since
$Z_{I}=\bigcap_{f\in I}\{x\in\mathscr{B}\mid x(f)=0\},$
the set $Z_{I}$ is closed in the set of points $X_{\mathscr{B}}$. We know that for any map of $C^{\infty}$-rings $\varphi:\mathscr{B}\to\mathscr{A}$ the induced map $f(\varphi):X_{\mathscr{A}}\to X_{\mathscr{B}}$, $f(\varphi)x=x\circ\varphi$, is continuous.
Let $J:=\ker\varpi$. By the first isomorphism theorem
$f(\varpi)\big{(}X_{\mathscr{A}}\big{)}=\{y\in X_{\mathscr{B}}\mid y|_{J}=0\},$
which is $Z_{J}$. Thus $f(\varpi)$ maps $X_{\mathscr{A}}$ continuously onto $Z_{J}$. Since $\varpi$ is onto, $f(\varpi)$ is injective. To prove that $f(\varpi):X_{\mathscr{A}}\to Z_{J}$ is a homeomorphism, it’s enough to prove that $f(\varpi)$ is an open map. A basic open set in $X_{\mathscr{A}}$ is of the form $\{a\not=0\}:=\{x\in X_{\mathscr{A}}\mid x(a)\not=0\}$. Since $\varpi$ is onto, there is $b\in\mathscr{B}$ such that $\varpi(b)=a$. Then
$x(\varpi(b))\not=0\quad\Leftrightarrow\quad(x\circ\varpi)(b)\not=0.$
Consequently
$f(\varpi)\,(\{a\not=0\})=\{y\in X_{\mathscr{B}}\mid y=f(\varpi)x\,\textrm{ and}\,x(a)\not=0\}=Z_{J}\cap\{b\not=0\},$
and we are done. ∎
A.iii. The spectrum of a $C^{\infty}$-ring $\operatorname{Spec}(C^{\infty}(M)^{G})$ of invariant functions.
We now turn to study $\operatorname{Spec}(\mathscr{A})$ where $\mathscr{A}$ is a ring of invariant functions $C^{\infty}(M)^{G}$ for an action of a Lie group $G$ on a manifold $M$.
Lemma A.6.
Let $G\times M\to M$ be an action of a Lie group $G$ on a manifold $M$. Then the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ of the invariant functions is point determined.
Proof.
For a point $p\in M$ we have a homomorphism
$\operatorname{\mathsf{ev}}_{p}:C^{\infty}(M)^{G}\to\mathbb{R},\quad\operatorname{\mathsf{ev}}_{p}(f):=f(p)$
If $f\in C^{\infty}(M)^{G}$ and $x(f)=0$ for all $x\in\operatorname{Hom}(C^{\infty}(M)^{G},\mathbb{R})$, then $f(p)=0$ for all $p\in M$, and then $f=0$. ∎
Let $G\times M\to M$ be an action. If two points $p,p^{\prime}\in M$ lie on the same $G$-orbit than for any invariant function $f\in C^{\infty}(M)$
$\operatorname{\mathsf{ev}}_{p}(f)\equiv f(p)=f(p^{\prime})\equiv\operatorname{\mathsf{ev}}_{p^{\prime}}.$
Hence the evaluation map
$\operatorname{\mathsf{ev}}:M\to X_{C^{\infty}(M)^{G}},\quad p\mapsto\operatorname{\mathsf{ev}}_{p}$
descends to a well-defined map
(A.5) $\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M)^{G}},\quad\overline{\operatorname{\mathsf{ev}}}(G\cdot p)=\operatorname{\mathsf{ev}}_{p}.$
In general there is no reason for the map $\overline{\operatorname{\mathsf{ev}}}$ to be a bijection. If invariant functions separate orbits (which happens when the action is proper) then $\overline{\operatorname{\mathsf{ev}}}$ is injective. In general it is not. For instance consider the action of the reals $\mathbb{R}$ on the torus $\mathbb{T}^{2}$ by irrational flow (cf. Subsection 4.i). Then the ring of invariants $C^{\infty}(\mathbb{T}^{2})^{\mathbb{R}}$ consists of constant functions. Consequently it has exactly one $\mathbb{R}$ point and the map $\overline{\operatorname{\mathsf{ev}}}:\mathbb{T}^{2}/\mathbb{R}\to X_{C^{\infty}(\mathbb{T}^{2})^{\mathbb{R}}}$ is not injective. We now argue that for proper actions the map $\overline{\operatorname{\mathsf{ev}}}$ is a homeomorphism. We do it in a series of lemmas.
Lemma A.7.
Let $M$ be a topological space and $\mathscr{A}\subset C^{0}(M)$ a $C^{\infty}$-subring. Suppose $\mathscr{A}$ contains a proper function. Then any unital $\mathbb{R}$-algebra map $\varphi:\mathscr{A}\to\mathbb{R}$ is an evaluation at a point $p\in M$:
$\varphi(g)=g(p)\quad\text{ for all $g\in\mathscr{A}$}.$
Consequently, all unital $\mathbb{R}$-algebra maps from $\mathscr{A}$ to $\mathbb{R}$ are morphisms of $C^{\infty}$-rings and the map
$M\to\operatorname{Hom}(\mathscr{A},\mathbb{R}),\quad x\mapsto\operatorname{\mathsf{ev}}_{x}$
is surjective. Here as before $\operatorname{\mathsf{ev}}_{x}(g):=g(x)$.
Proof.
 [^1] I am greatful to Yael Karshon for fixing a gap in an earlier version of the proof.
We use the same notation for a real number and for the corresponding constant function in $C^{0}(M)$. Since $\mathscr{A}$ is a $C^{\infty}$-subring of $C^{0}(M)$, it contains these constant functions.
Let $\varphi:\mathscr{A}\to\mathbb{R}$ be a unital homomorphism of $\mathbb{R}$-algebras. For $g\in\mathscr{A}$ set
$U_{g}:=\{p\in M\mid\varphi(p)\neq g(p)\},$
We now argue by contradiction. Assume that for every $p\in M$ there exists $g\in\mathscr{A}$ such that $\varphi(p)\neq g(p)$. Then $\{U_{g}\}_{g\in\mathscr{A}}$ is an open cover of $M$. Let $f\in\mathscr{A}$ be a proper function, which exists by our assumption. Let
$r:=\varphi(f).$
Since $f$ is proper, the level set $\{f=r\}$ is compact. Therefore there exists $n>0$ and $g_{1},\ldots,g_{n}\in\mathscr{A}$ such that
$\{f=r\}\subseteq U_{g_{1}}\cup\ldots\cup U_{g_{n}}.$
Then the function
$h(\cdot):=\big{(}f(\cdot)-r\big{)}^{2}+\big{(}g_{1}(\cdot)-\varphi(g_{1})\big{)}^{2}+\ldots+\big{(}g_{n}(\cdot)-\varphi(g_{n})\big{)}^{2}\in\mathscr{A}$
is strictly positive. Since $\varphi$ is a homomorphism of $\mathbb{R}$-algebras, $\varphi(h)=0$. Since $h\geq(f(\cdot)-r)^{2}$ and since $(f(\cdot)-r)^{2}$ is proper, the function $h$ is proper as well. Since $h$ is proper and since $\mathbb{R}$ is locally compact Hausdorff, $h$ is a closed map. It follows, since $h$ is positive, that there is $\varepsilon>0$ so that $h(m)>\varepsilon$ for all $m\in M$. There exists a smooth function $k\in C^{\infty}(\mathbb{R})$ such that
$k(y)=\frac{1}{y}\quad\text{ for all \ }y>\varepsilon.$
Note that $k\circ h\in\mathscr{A}$ since $\mathscr{A}$ is a $C^{\infty}$-subring of $C^{0}(M)$ and the $C^{\infty}$-ring operations on $C^{0}(M)$ are given by composition. Since $h\cdot(k\circ h)\equiv 1$ and $\varphi$ is an $\mathbb{R}$-algebra morphism, $\varphi(h)\cdot\varphi(k\circ h)=\varphi(1)$. But $\varphi(h)=0$, so $\varphi(1)=0$, which contradicts that $\varphi$ is unital. ∎
Remark A.8.
Let $G\times M\to M$ be an action of a Lie group $G$ on a manifold $M$. Let $q:M\to M/G$ denote the orbit/quotient map. By the universal property of $q$ for any $G$-invariant function $f:M\to\mathbb{R}$ there exists a unique continuous function $\overline{f}:M/G\to\mathbb{R}$ with $f=\overline{f}\circ q$. We thus get an injective map
(A.6) $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G),\qquad f\mapsto\overline{f}$
of $C^{\infty}$-rings. We define
(A.7) $C^{\infty}(M/G):=\Upsilon(C^{\infty}(M)^{G})\qquad$
Then
$\Upsilon:C^{\infty}(M)^{G}\to C^{\infty}(M/G)$
is an isomorphism of $C^{\infty}$-rings.
Lemma A.9.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$. Then the $C^{\infty}$-subring $C^{\infty}(M/G)$ of $C^{0}(M/G)$ defined in Remark A.8 contains a proper function. Consequently the map
$\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M/G)},\quad\overline{\operatorname{\mathsf{ev}}}(G\cdot p)(f)=f(p)\textrm{ for all }f\in C^{\infty}(M)^{G}$
is a bijection.
Proof.
As was mentioned before, since functions invariant under proper actions separate orbits, the map $\overline{\operatorname{\mathsf{ev}}}$ is injective.
In the course of the proof of Theorem 4.3.1 in [Pa] Palais constructs a sequence of invariant functions $\{f_{n}\}_{n\in\mathbb{N}}\subset C^{\infty}(M)^{G}$ with the property that their supports form a locally finite cover of $M$, that $f_{n}(M)\subset[0,1]$ for all $n$, $\sum f_{n}=1$ and that sets $q(\operatorname{supp}(f_{n}))\subset M/G$ are compact (here as before $q:M\to M/G$ is the quotient map). Let $\overline{f}_{n}$ denote the image of $f_{n}$ in $C^{0}(M/G)$ under the map $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G)$ (cf. Remark A.8). Then the function
$h=\sum_{n=1}^{\infty}n\overline{f}_{n})$
is the desired proper function (cf. proof of Proposition 2.28 in [Lee]). ∎
Lemma A.10.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$. Then the space of points $X_{C^{\infty}(M)^{G}}$ of the $C^{\infty}$-ring of invariant functions is homeomorphic to the orbit space $M/G$.
Proof.
As we observed in Remark A.8 the map $\Upsilon:C^{\infty}(M)^{G}\to C^{\infty}(M/G)$ is an isomorphism of $C^{\infty}$-rings, hence $X_{C^{\infty}(M)^{G}}$ and $X_{C^{\infty}(M/G}$ are homeomorphic. By Lemma A.9 the evaluation map
$\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M)^{G}},\quad\overline{\operatorname{\mathsf{ev}}}(G\cdot p)(f)=f(p)\textrm{ for all }f\in C^{\infty}(M)^{G}$
is a bijection. We need to show that $\operatorname{\mathsf{ev}}$ is a homeomorphism. By Lemma A.11 below the topology on $M/G$ is generated by the sets of the form $\{h\not=0\}:=\{p\in M/G\mid h(p)\not=0\}$, $h\in C^{\infty}(M/G)$. On the other hand the topology on $X_{C^{\infty}(M/G)}$ is generated by the sets of the form
$\{f\not=0\}:=\{x\in X_{C^{\infty}(M/G)}\mid x(f)\not=0\},$
$f\in C^{\infty}(M/G)$. Now observe that
$(\operatorname{\mathsf{ev}})^{-1}(\{x\in X_{C^{\infty}(M/G)}\mid x(f)\not=0\})=\{p\in M/G\mid f(p)\not=0\}.$
Hence $\operatorname{\mathsf{ev}}$ is a homeomorphism. ∎
Lemma A.11.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$, $q:M\to M/G$ the quotient map and $C^{\infty}(M/G)\subset C^{0}(M/G)$ defined by (A.7). The quotient topology on $M/G$ is the smallest topology making all the functions in $C^{\infty}(M/G)$ continuous. That is, the quotient topology is generated by the sets of the form $\{h\not=0\}$ for all $h\in C^{\infty}(M/G)$.
Proof.
By [KL2][Lemma 2.15] it is enough to show that for any point $p\in M/G$ and for any closed set $C\subset M/G$ with $p\not\in C$ there is a function $f\in C^{\infty}(M/G)$ with $\operatorname{supp}(f)\cap C=\varnothing$ and $f$ identically 1 near $p$.
[…]
Lemma A.12.
Let $G\times M\to M$ be a proper action of a Lie group $G$ on a manifold $M$ and let $q:M\to M/G$ denote the quotient map as before.. Then the affine $C^{\infty}$-scheme $\operatorname{Spec}(C^{\infty}(M^{G}))$ of the ring of invariant functions is isomorphic to $(M/G,C^{\infty}_{M/G})$ where
$C^{\infty}_{M/G}(U)=\{f:U\to\mathbb{R}\mid f\circ q\in C^{\infty}(q^{-1}(U))\}$
for all open sets $U\subset M/G$.
Proof.
By Lemma A.10 the space of points $X_{C^{\infty}(M)^{G}}$ is homeomorphic to the orbit space $M/G$ by way of the map
$\overline{\operatorname{\mathsf{ev}}}:M/G\to X_{C^{\infty}(M)^{G}},\quad\left(\overline{\operatorname{\mathsf{ev}}}(G\cdot p)\right)(f)=f(p).$
Consequently
(A.8) $(\overline{\operatorname{\mathsf{ev}}})^{*}:C^{0}(X_{C^{\infty}(M)^{G}})\to C^{0}(M/G),\quad h\mapsto h\circ\overline{\operatorname{\mathsf{ev}}}$
is an isomorphism of $C^{\infty}$-rings.
By Lemma A.6 the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is point determined, hence can be identified with a $C^{\infty}$-subring of $C^{0}(X_{C^{\infty}(M)^{G}})$ by way of
(A.9) $(\,\,)_{*}:C^{\infty}(M)^{G}\to C^{0}(X_{C^{\infty}(M)^{G}}),\quad f\mapsto f_{*}$
where $f_{*}(x):=x(f)$ for all $f\in C^{\infty}(M)^{G}$ and all $x\in X_{C^{\infty}(M)^{G}}$. Since for any $f\in C^{\infty}(M)^{G}$
$((\overline{\operatorname{\mathsf{ev}}})^{*}f_{*})(G\cdot p)=\left(\operatorname{\mathsf{ev}}(G\cdot p)\right)(f)=f(p)$
the composite of (A.9) and (A.8) is the map $\Upsilon:C^{\infty}(M)^{G}\to C^{0}(M/G)$ that sends the invariant function $f\in C^{\infty}(M)^{G}$ to the unique function $\overline{f}:M/G\to\mathbb{R}$ with $\overline{f}\circ q=f$ (cf. Lemma A.8). Lemma A.3 now implies the desired result. ∎
A.iv. $\operatorname{Spec}(C^{\infty}(M)^{G})$ when the action $G\times M\to M$ is proper and $C^{\infty}(M)^{G}$ is finitely generated.
We now put together Subsections A.i and A.iii. Suppose now $G\times M\to M$ is a proper action and the $C^{\infty}$-ring of invariants $C^{\infty}(M)^{G}$ is finitely generated (hence there is a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to C^{\infty}(M)^{G}$ for some $N$). Then, one one hand,
$\operatorname{Spec}(C^{\infty}(M)^{G})\simeq(Z,C^{\infty}_{Z})$
where as before
$Z=\{p\in\mathbb{R}^{N}\mid f(p)=0\textrm{ for all }f\in\ker\varpi\}$
is the zero set of the kernel of $\varpi$ and the sheaf $C^{\infty}_{Z}$ is given by
$C^{\infty}_{Z}(U)=C^{\infty}(\widetilde{U})|_{U}$
for any $\widetilde{U}\in\mathsf{Open}(\mathbb{R}^{n})$ with $\widetilde{U}\cap Z=U$. On the other hand
$\operatorname{Spec}(C^{\infty}(M)^{G})\simeq(M/G,C^{\infty}_{M/G}).$
By composing the two isomorphisms of local $C^{\infty}$-ringed spaces we get an isomorphism
$\underline{\varphi}=(\varphi,\varphi_{\#}):(M/G,C^{\infty}_{M/G})\to(Z,C^{\infty}_{Z}).$
We now write down $\underline{\varphi}$ explicitly. Let $x_{i}:\mathbb{R}^{N}\to\mathbb{R}$, $1\leq i\leq N$, denote the coordinate functions. Then their images $f_{i}=\varpi(x_{i})$, $1\leq i\leq N$, generate $C^{\infty}(M)^{G}$. Let
$F=(f_{1},\ldots,f_{N}):M\to\mathbb{R}^{N}.$
Then
$g\circ F=\varpi(g)$
for all $g\in C^{\infty}(\mathbb{R}^{N})$, i.e., $\varpi=F^{*}$. The closed embedding
$\operatorname{Spec}(\varpi):(X_{C^{\infty}(M)^{G}},\mathscr{O}_{C^{\infty}(M)^{G}})\simeq(M/G,C^{\infty}_{M/G}\to\operatorname{Spec}(C^{\infty}(\mathbb{R}^{N})=(\mathbb{R}^{N},C^{\infty}_{\mathbb{R}^{N}})$
is then $\operatorname{Spec}(\varpi)=(\overline{F},F_{\#})$ where $\overline{F}:M/G\to\mathbb{R}^{N}$ is the continuous map induced by $F:M\to\mathbb{R}^{N}$ (i.e., $\overline{F}(G\cdot p)=F(p)$) and $F_{\#}:C^{\infty}_{\mathbb{R}^{N}}\to C^{\infty}_{M/G}$ is
$F_{\#,W}:C^{\infty}(W)\to C^{\infty}_{M/G}((\overline{F})^{-1}(W)0\simeq C^{\infty}(F^{-1}(W)^{G},\quad h\mapsto h\circ F$
for all open sets $W\subset\mathbb{R}^{N}$. Then
$\overline{F}(M/G)=F(M)=Z$
and $\ker(F_{\#})$ is the ideal sheaf of $Z$. Thus $\operatorname{Spec}(\varpi)$ factors through
$(\varphi,\varphi_{\#}):(M/G,C^{\infty}_{M/G})\to(Z,C^{\infty}_{Z})$
where $\varphi(G\cdot p)=F(p)$ and
$\varphi_{\#,W}:C^{\infty}_{Z}(W)\to C^{\infty}(F^{-1}(W)^{G},\quad\varphi_{\#,W}(h)=h\circ F.$
Appendix B A remark on the Cushman-Weinstein conjecture
Suppose $M$ is a symplectic manifold with a proper Hamiltonian action of a Lie group $G$ and an associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Suppose further that the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is finitely generated. Then the $C^{\infty}$-ring of invariants $C^{\infty}(M)^{G}$ is Poisson and consequently $\operatorname{Spec}(C^{\infty}(M)^{G})\simeq(M/G,C^{\infty}_{M/G})$ is a Poisson $C^{\infty}$-scheme. Moreover as we saw in A.iv a choice of a surjective map $\varpi:C^{\infty}(\mathbb{R}^{N})\to C^{\infty}(M)^{G}$ (i.e., a choice of generators of $C^{\infty}(M)^{G}$) gives rise to a closed embedding
$\operatorname{Spec}(\varpi):(M/G,C^{\infty}_{M/G})\to(\mathbb{R}^{N},C^{\infty}_{\mathbb{R}^{N}}),$
which induces an isomorphism
$(M/G,C^{\infty}_{M/G})\to(Z,C^{\infty}_{Z})$
for an appropriate closed subset $Z$ of $\mathbb{R}^{N}$. In particular the $C^{\infty}$-ring
$C^{\infty}_{Z}(Z)=C^{\infty}(Z)=C^{\infty}(\mathbb{R}^{N})|_{Z}$
carries a Poisson bracket. In 1980s Cushman [SL] and, independently, Weinstein [Eg1, Eg2] conjectured that the Poisson bracket on $C^{\infty}(Z)$ extends to a Poisson bracket on $C^{\infty}(\mathbb{R}^{N})$. Egilsson [Eg1, Eg2] showed that the conjecture is false as stated. Further work in this direction was done by Davis [Da] and McMillan [Mc]. In this section we make the following observation.
Lemma B.1.
[…]
$q(\langle df\wedge dg,\Lambda\rangle)=\{q(f),q(g)\}$
for all $f,g\in C^{\infty}(\mathbb{R}^{N})$.
Proof.
Recall that the $C^{\infty}$-ring $C^{\infty}(\mathbb{R}^{N})$ is freely generated by the coordinate functions $x_{1},\ldots,x_{N}:\mathbb{R}^{N}\to\mathbb{R}$. More concretely, for any $f\in C^{\infty}(\mathbb{R}^{N})$ we have
$f=(f)_{C^{\infty}(\mathbb{R}^{N})}(x_{1},\ldots,x_{N}).$
Since by assumption $q:C^{\infty}(\mathbb{R}^{N})\to\mathscr{B}$ is surjective, for all indicies $i$ and $j$ there exist $c_{ij}\in C^{\infty}(\mathbb{R}^{N})$ so that $q(c_{ij})=\{q(x_{i}),q(x_{j})\}$. Define the bivector field $\Lambda$ by
$\Lambda=\frac{1}{2}\sum_{i,j}c_{ij}\partial_{i}\wedge\partial_{j}.$
Then
$\langle df\wedge dg,\Lambda\rangle=\sum_{i,j}c_{ij}\partial_{i}f\,\partial_{j}g.$
Consequently
$\begin{split}q(\langle df\wedge dg,\Lambda\rangle)&=\sum_{ij}q(c_{ij})q(\partial_{i}f)\,q(\partial_{j}g)\\ &=\sum_{ij}(\partial_{i}f)_{\mathscr{B}}(q(x_{1}),\ldots,q(x_{N})(\partial_{j}g)_{\mathscr{B}}(q(x_{1}),\ldots,q(x_{N})\{q(x_{i}),q(x_{j})\}\\ &=\{f_{\mathscr{B}}(q(x_{1}),\ldots,q(x_{N}),g_{\mathscr{B}}(q(x_{1}),\ldots,q(x_{N})\}=\{q(f),q(g)\}\end{split}$
for all $f,g\in C^{\infty}(\mathbb{R}^{N})$.
It follows that if $(M,\omega,\mu:M\to\mathfrak{g}^{\vee})$ is a Hamiltonian $G$-space, the action of $G$ on $M$ is proper and the ring of invariants $C^{\infty}(M)^{G}$ is finitely generated with generators $f_{1},\ldots f_{N}$ then
$F:M\to\mathbb{R}^{N},\quad F(m)=(f_{1}(m),\ldots f_{N}(m))$
induces an embedding $\overline{F}:M/G\to\mathbb{R}^{N}$ so that $Z=\overline{F}(M/G)$ is closed in $\mathbb{R}^{N}$ and ${\overline{F}}^{*}:C^{\infty}(Z)\to C^{\infty}(M/G)$ is an isomorphism of $C^{\infty}$-rings. Consequently $C^{\infty}(Z)=C^{\infty}(\mathbb{R}^{N})|_{Z}$ is a Poisson $C^{\infty}$-ring, and the Poisson bracket on $C^{\infty}(Z)$ extends (non-uniquely) to a bivector field on $\mathbb{R}^{N}$.